### v1

Abstract
Objective: This study proposes a Natural Language Processing pipeline to evaluate prescription appropriateness in Italian referrals, where reasons for prescriptions are recorded only as free text, complicating automated comparisons with guidelines. The pipeline aims to derive, for the first time, a comprehensive summary of the reasons behind these referrals and a quantification of their appropriateness. While demonstrated in a specific case study, the approach is designed to generalize to other types of examinations.
Methods: Leveraging embeddings from a transformer-based model, the proposed approach clusters referral texts, maps clusters to labels, and aligns these labels with existing guidelines. We present a case study on a dataset of 496,971 referrals, consisting of all referrals for venous echocolordopplers of the lower limbs between 2019 and 2021 in the Lombardy Region. A sample of 1,000 referrals was manually annotated to validate the results.
Results: The pipeline exhibited high performance for referrals’ reasons (Prec=92.43%, Rec=83.28%) and excellent results for referrals’ appropriateness (Prec=93.58%, Rec=91.52%) on the annotated subset. Analysis of the entire dataset identified clusters matching guideline-defined reasons - both appropriate and inappropriate - as well as clusters not addressed in the guidelines. Overall, 34.32% of referrals were marked as appropriate, 34.07% inappropriate, 14.37% likely inappropriate, and 17.24% could not be mapped to guidelines.
Conclusions: The proposed pipeline effectively assessed prescription appropriateness across a large dataset, serving as a valuable tool for health authorities. Findings have informed the Lombardy Region’s efforts to strengthen recommendations and reduce the burden of inappropriate referrals.
keywords:
Natural language processing, Referrals, Prescription appropriateness, Public healthcare system, Quality of healthcare
1 Background and significance
In Italy, the National Health System (NHS) is the primary provider of healthcare services, offering universal coverage, with decentralized administration at the regional level [1]. Despite a strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, leading many citizens to turn to private providers [2].
One significant contributor to these delays is the issue of prescription appropriateness, a concern shared by many advanced healthcare systems [3, 4]. Inappropriate prescriptions have a threefold negative impact on the healthcare system: waste of resources, increased waiting times, and reduced effectiveness of the care provided to citizens, who face delays in obtaining accurate diagnoses [5]. In the Italian NHS, referrals from general practitioners (GPs) or specialized physicians are necessary to book examinations. Assessing prescription appropriateness requires understanding the reasons underlying these referrals.
When Italian physicians fill out a referral, they have to use a standardized form, providing the reason for the prescription in a specific free-text field called Clinical Question. The unstructured nature of this field, coupled with the lack of standardized alternatives, makes it challenging to analyze. Addressing this issue requires Natural Language Processing (NLP) techniques.
The number of studies analyzing texts of Italian referrals is very limited. Venturelli et al. [6] used proprietary commercial software (Clinika VAP [7]), for which the NLP models or algorithms are not disclosed. This software provides an appropriate/inappropriate label for each referral, enabling the quantification of inappropriate referrals. However, it does not provide a clustered analysis of the referrals, with a quantification of the different prescription reasons. To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [8], which focused on extracting and analyzing follow-up waiting times from clinical questions to investigate timely access to care.
While several studies have evaluated prescription appropriateness for medications within the Italian NHS using structured data or manual text analysis (e.g., [9, 10, 11]), research on the appropriateness of examinations is more limited [12, 13] and has not incorporated natural language processing.
A few studies developed decision-support systems for guideline adherence, both in primary care [14, 15] and hospital settings [16, 17]. These tools assist physicians through closed-form questions, possibly complemented by keyword matching on the text they write. While effective in specific scenarios [18, 19], such tools are not suited for retrospective analyses of their performance or for quantifying the prevalence of appropriate or inappropriate referrals.
Beyond the Italian context, Haddad et al. [20] analyzed Russian referrals’ texts for prescription appropriateness using labelled data; Sagheb et al. [21] designed a rule-based algorithm to verify guidelines adherence for asthma patients from EHR documents; similarly in [22] a rule-based approach was applied to assess guideline adherence for gout patients. Dimitrov et al. [23] conducted a nationwide evaluation of breast cancer guideline adherence in EHRs using proprietary commercial software (Danny platform [24]), which incorporated rules and manual reviews. Villeana et al. [25], analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
These studies, however, are often disease-specific or reliant on labelled data or proprietary software, limiting their broader applicability.
2 Objective
This study proposes an NLP-based pipeline to assess the appropriateness of Italian referrals, developing a model designed for unsupervised analysis across different types of medical examinations.
In 2015, the Lombardy Region, the most populated Italian region with approximately 10 million inhabitants - akin to a small state - issued two resolutions (DGRs) focusing on prescription appropriateness in angiology [26] and gastroenterology [27], two fields where the literature reports high rates of inappropriate prescriptions, both in Italy and internationally [28, 29, 30, 31, 32]. These resolutions established guidelines specifying appropriate and inappropriate clinical questions for a variety of examinations.
Building on these efforts, we analyzed examinations covered by these resolutions, contributing to the issuance of a new resolution in December 2023 that reinforced the previous guidelines [33]. This paper focuses on a case study of one of these examinations: the venous echocolordoppler of the lower limbs. In particular, we analyzed a dataset consisting of all referrals prescribed in the Lombardy Region for this exam between 2019 and 2021, for a total of 496,971 referrals, encompassing 229,509 unique clinical questions.
The analysis presented in this study aims to verify the effectiveness of the 2015 resolutions in enforcing prescriptive appropriateness, providing for the first time an overview of the main reasons why these examinations are prescribed, with an approximate quantification of the number of referrals for each disease/symptom and their alignment on existing guidelines, where possible.
Given the absence of labelled data and the need to uncover reasons not addressed in guidelines, we framed this as a clustering problem. In particular, our pipeline clusters clinical questions by exploiting the embeddings produced by a fine-tuned BERT model [34]. These clusters are summarized and, where possible, mapped to guidelines to determine whether they are appropriate, inappropriate, or not covered by the guidelines.
A subset of 1,000 distinct clinical questions was manually annotated to validate the pipeline. Results are reported for this validation subset to verify the correctness of the pipeline output and for the full dataset to analyze prescription appropriateness at the regional level, including stratifications by year (pre-, during, and post-COVID pandemic), geographical area, and type of prescribing physician. Additional results for another type of examination in the gastroenterological domain are included in Appendix D.
The main contributions of this study are the following:
  1. 1.
    The first open-source, fully unsupervised pipeline for assessing prescription appropriateness from Italian referrals, applicable to any type of examination.
  2. 2.
    The first region-wide analysis in Italy of referrals’ reasons and their appropriateness.
3 Material and Methods
3.1 Data
3.1.1 Referrals dataset
The dataset analyzed in this study is extracted from all referrals for venous echocolordoppler of the lower limbs prescribed in the Lombardy Region between 2019 and 2021 (496,971 referrals). Each referral includes the clinical question, the year, the Local Health Authority (LHA) of the prescribing physician and the type of prescribing physician. Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs represent the eight healthcare administrative areas within the Lombardy Region. Figure 1 illustrates the distribution of referrals among these groups.
The analysis focuses on the clinical question, a free-text field typically containing one or few sentences. These texts exhibit frequent spelling and typing errors, heavy use of abbreviations, inconsistent punctuation, and non-standard capitalization. The dataset includes 229,509 distinct clinical questions, but only 15% appear more than once. The median length is 71 characters (IQR=39). Only 189 (0.08%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
The clinical questions should not contain personal information about patients or physicians. A preliminary check was performed by the Lombardy Region IT provider, masking any suspect personal information before data extraction.
3.1.2 Manual annotations
To validate the identified clusters, a randomly selected subset of 1,000 distinct clinical questions was manually labelled. The labels correspond to categories representing diseases or symptoms that prompted the referral. During an initial review of this subset, ambiguous cases were identified, and an angiologist was consulted to solve doubts and align clusters with prescription guidelines. Annotation guidelines were subsequently refined, and the entire subset was annotated accordingly. Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled dataset was error-free. More details on the annotations are provided in Appendix A.
3.1.3 Prescription guidelines
The guidelines used as a reference to assess referral appropriateness are extracted from [26]. In particular, we extracted the appropriate and inappropriate diseases and symptoms for the venous echocolordoppler of the lower limbs. An English translation is reported in Figure 2.
Figure 1: Referrals distributions across years (A), type of prescribing physicians (B) and LHA (C)
Figure 2: Guidelines for referrals for venous echocolordoppler of the lower limbs in the Lombardy Region. Translated from the Appendix 3 of [26].
3.2 Pipeline for prescription appropriateness
In this section, we present our pipeline to evaluate prescription appropriateness. The process can be divided into two main steps:
  1. 1.
    Identify the reason for the referral: extract the central semantic meaning from the clinical question.
  2. 2.
    Verify appropriateness: determine if the identified reason aligns with established guidelines.
The absence of labels for this type of data prevented us from adopting a supervised approach. A rule-based approach was excluded since we aim to build a pipeline that can generalize to any examination, and it would not be feasible to develop tailored rules for each type of examination. Moreover, we are interested in understanding the reasons for these referrals, which might include additional and/or more detailed categories than those mentioned in the guidelines. Because of this, the problem was addressed as a clustering problem.
Figure 3 depicts blocks composing our pipeline, which are detailed below.
Figure 3: Schema of the pipeline for assessing referrals reasons and appropriateness
3.2.1 Pre-processing
Physicians write referrals during patients’ examinations, with limited time available. This leads them to write clinical questions that are far from well-readable and correct texts. The grammar of the sentences is often poor, there are frequent typing errors, and the punctuation is either not used or used incorrectly. The use of uppercase letters does not have any particular meaning, and the use of abbreviations is very frequent. Even in the single type of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms. In the manually labelled subset, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
To address these issues, suitable pre-processing steps were applied:
  1. 1.
    lowercasing
  2. 2.
    punctuation removal
  3. 3.
    abbreviation expansion
  4. 4.
    (partial) typos correction
We removed all punctuation and lowercase sentences due to the inconsistent use of these elements.
[…]
Examples of pre-processing effects are reported in Appendix C.
3.2.2 Text embedding
To appropriately cluster the clinical questions, it is necessary to produce a representation capturing the central semantics of these texts without getting confused by their widely present noise. Because of this, vocabulary-based representations, such as TF-IDF [35], cannot be expected to work well for this task. The main alternatives are static and contextual neural network-based embeddings.
Static word embeddings, such as Word2Vec [36], can capture the semantic similarities between words with similar meanings (e.g., vasculopathy - arteriopathy, phlebitis - thrombophlebitis) differently from vocabulary-based representations. Nevertheless, they might have more difficulty distinguishing between relevant and irrelevant tokens in our sentences since they do not consider the context in which tokens appear.
On the contrary, context-based embeddings produced by transformer models consider the surrounding context while embedding each token, and have often shown superior performances in the recent clinical NLP literature [37, 38, 39, 40]. The problem is particularly challenging in our case due to the limited number of embedding models available for the Italian language, especially for the clinical domain.
To this purpose, the same model used in [41] was adopted to cluster sentences extracted from Italian referrals, and we verified that even with clinical questions, its performances are superior to other models. This model consists of a fine-tuned version of the last Italian version of BERT, Umberto [42], on a dataset of publicly available medical documents in Italian and other languages (translated into Italian) [43].
3.2.3 Clustering
HDBscan was adopted to cluster the embeddings since it does not require eliciting in advance the number of clusters and it allows a not complete assignment of items to clusters. HDBscan requires defining the minimum size of the clusters to identify. We verified that the best results are achieved when this parameter is progressively reduced while iterating the clustering on the previously unclustered data. This approach reduced the number of unclustered points, addressing the varying cluster sizes in our dataset.
An additional issue in the clustering procedure was related to the size of our dataset. On a machine with 32 GB of RAM, it was not possible to run the clustering algorithm on the whole dataset. Because of this, we had to divide it into chunks of 30,000 clinical questions and cluster them separately. The subsequent summarization and mapping steps merged the clusters found in different chunks and corresponding to the same diseases/symptoms.
3.2.4 Cluster summarization
Summarizing the identified clusters is necessary to understand the symptoms or diseases corresponding to each cluster. To automatize this step, we extracted keywords from each cluster. The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold until at least two keywords were found. Among these, we excluded those with high frequency in at least half of the clusters. This allowed us to exclude not only stop words but also frequent words for the specific exam (e.g., leg, lower limbs, right, left in our case study).
3.2.5 Mapping clusters to guidelines
After summarization, the keyword representation of the clusters was used to map them to the guidelines. The process involved:
  1. 1.
    Computing embeddings for each clinical question described in the guidelines.
  2. 2.
    Generating an embedding for each cluster based on its keyword summary, enabling direct comparison with the guideline embeddings.
  3. 3.
    Comparing cluster embeddings with guideline embeddings using cosine similarity. Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50absent0.50>0.50> 0.50.
3.3 Experiments and evaluation
3.3.1 Pipeline evaluation
To validate the pipeline, we compared its output with labels from the manually annotated subset. The pipeline was executed on the entire dataset, and results for the manually labelled portion were extracted. This approach ensures fairness, as running the pipeline solely on the labelled subset (1,000 records) would limit its ability to form meaningful clusters due to the restricted data size.
Manual annotations identified 49 distinct labels, with only 13 occurring more than 10 times. For evaluation purposes, these 13 labels were retained for comparison, while the remaining labels were grouped into the other category, which we expect to align with unclustered elements.
The results were evaluated in terms of precision, recall, and f1-score. Metrics were calculated at two levels: at aggregate level to evaluate the overall pipeline performance across all labels, and at per-label level to evaluate performance for individual labels.
Additionally, we weighted clinical questions by their corresponding referral counts, incorporating the dataset’s real-world distribution into the analysis.
A second type of evaluation is performed to evaluate the quality of prescription appropriateness labels. In this case, we grouped the 13 labels into four categories: appropriate, inappropriate, likely inappropriate and undetermined. In particular, we marked as likely inappropriate some clusters for which the indication was not completely clear (see Results section for details on this), and as undetermined those that were not covered by the guidelines.
In both cases, we compared our pipeline with other types of text representations/clustering algorithms. We considered clustering results obtained by embeddings from the original Umberto model, by embeddings from an Italian Word2Vec model [44], by TF-IDF representations and by the LDA algorithm. See Appendix B for details.
3.3.2 Analysis of the entire dataset
The pipeline is applied to the entire dataset, with results that provide a complete quantification of referrals for each identified category, together with their appropriateness. Moreover, we present stratified results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs specialists).
4 Results
Table 1: Results of clusters mapped to the labels defined by annotations on the manually labelled test set, computed considering clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
Umberto-E3C-Pipeline 86.81 77.32 79.48 92.43 83.28 86.49
Umberto-Base 82.11 73.62 77.21 87.45 77.81 81.89
Word2Vec 76.21 67.55 71.24 81.67 72.35 76.73
TF-IDF 50.11 41.20 42.43 52.54 44.81 45.45
LDA 27.74 34.04 28.48 19.19 27.57 29.90
Table 2: Results of the cluster mapping for each category, on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF). Labels with support <10absent10<10< 10 were grouped with other. P=Precision, R=Recall, F1=F1-Score. P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
Category P-CQ R-CQ F1-CQ # CQ P-REF R-REF F1-REF # REF
Claudicatio 83.33 45.45 58.82 11 90.00 56.25 69.23 16
Pain 83.61 91.07 87.18 56 86.42 93.33 89.74 75
Edema 85.47 96.15 90.50 104 89.51 97.32 93.25 149
Venous insuff. 94.90 74.50 83.47 200 98.06 88.09 92.81 747
Paresthesis 100.00 81.25 89.66 16 100.00 88.00 93.62 25
Saphenectomy 100.00 92.59 96.15 27 100.00 93.94 96.88 33
(Trombo) Phlebitis 98.70 93.83 96.20 81 99.09 93.16 96.04 117
Thrombosis 96.24 94.04 95.13 218 96.99 94.77 95.87 306
Ulcer 95.83 92.00 93.88 25 97.44 92.68 95.00 41
Varices 41.43 89.23 56.59 65 82.47 98.89 89.83 628
Varices surgery 100.00 42.86 60.00 14 100.00 46.67 63.64 15
Vasculopathy 82.14 44.23 57.50 52 95.33 68.92 80.00 148
Other 65.37 67.94 65.68 131 62.56 70.59 66.33 187
4.1 Pipeline evaluation
Table 1 presents the results of the evaluation performed on the manually annotated subset. This evaluation compares the labels assigned to the clustered data with the ones defined by manual annotation. Clusters not mapped to any label or unclustered clinical questions were assigned to the other category.
The pipeline demonstrated superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations. The metrics were computed for both clinical questions (CQ) and referrals (REF). The referral-level scores were consistently higher due to the influence of clinical questions that occur more frequently, and that seem to be easier to cluster, being less detailed.
Figure 4: Confusion matrix of the clustering assignment to the labels on the manually annotated test set, computed considering the number of clinical questions (A) and the number of referrals (B).
[…]
The confusion matrices shown in Figure 4 allow for a more detailed analysis of misclassification errors. Errors in varices surgery are often due to misclassifications as varices, a strictly related cluster. Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other. Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we noticed that some presented both diagnoses, making this misclassification reasonable. For the remaining classes, misclassified samples fall mainly into the other group, affecting its precision.
Table 3: Results of clusters mapped to the appropriate/inappropriate/likely inappropriate/undetermined categories on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score. The values of all metrics are expressed as percentages.
P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
LDA 47.61 48.43 47.83 42.64 46.31 41.31
Table 3 shows the results achieved by grouping labels into the appropriate, inappropriate, likely inappropriate and undetermined categories. The results show higher aggregate performance scores compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories. This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging. A detailed list of all identified clusters, together with their mapping, is reported in Appendix B.
Similar results, including validation on a manually annotated subset, are reported in Appendix D for another type of examination.
4.2 Analysis of the entire dataset
Table 4: Summary of referrals’ clusters, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters more likely to be inappropriate, and light grey those not mentioned in the guidelines.
Cluster Label # REF %
Venous insufficiency 102,339 20,63
Thrombosis 98,821 19,92
Varices 71,269 14,37
Edema 61,822 12,46
(thrombo)phlebitis 49,537 9,99
Pain 21,131 4,26
Vasculopathy 11,244 2,27
Saphenectomy 8,290 1,67
Varices surgery 7,126 1,44
Diabetes 6,583 1,33
Ulcer 6,272 1,26
Paresthesias 4,856 0,98
Claudicatio 3,570 0,72
Dimer 1,601 0,32
Swelling 1,238 0,25
Respiratory insufficiency 1,060 0,21
Heavy legs 705 0,14
Ischaemia 617 0,12
Stent 612 0,12
Dermatitis 474 0,10
Hypercolesterolaemia 416 0,08
Pregnancy 225 0,05
Dialisis 222 0,04
Surgery 198 0,04
Smoker 193 0,04
Dyspnoea 171 0,03
Tingling 170 0,03
Cold legs 147 0,03
Hypoesthesia 80 0,02
Metabolic syndrome 48 0,01
Bleeding 29 0,01
Other 34,917 7,04
Figure 5: Examples of clustered clinical questions. English translations do not completely reflect the abbreviations and typos of the original Italian texts.
Table 4 extends the analysis on the entire dataset, highlighting the appropriateness labels for each cluster. It can be observed that the distribution of the reasons for the referrals is highly unbalanced, with the first five covering more than 75% of referrals. Varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to completely exclude, even for a human reader. Summarizing the table, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate. The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations. Notably, the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines.
Figure 5 provides some examples, along with their English translation, of clustered clinical questions.
Figure 6: Prescription appropriateness grouped by LHA (A), type of prescribing physicians (B) and year (C)
Figure 6 reports the level of appropriateness stratified by LHA, type of prescribing physicians and year, providing additional insights into the reasons for inappropriate prescriptions.
5 Discussion
In this study, we analyzed texts of Italian referrals for prescription appropriateness, with a case study on venous echocolordoppler of lower limbs. By introducing a novel approach that clusters referral texts and maps these clusters to existing guidelines, we demonstrated how it is possible to extract meaningful and actionable insights into the reasons behind such prescriptions.
Our work has several strengths. The analysis encompassed 496,971 referrals, covering an entire Italian region with 10 million inhabitants for three years. This large-scale dataset enabled a comprehensive evaluation of prescription practices. Validation was performed on a manually annotated subset, demonstrating that contextual embedding models significantly outperformed simpler clustering techniques in capturing the semantic nuances of referral texts. Analyzing the results in detail, we observed very high performances for most diseases/symptoms, with only two groups with lower recall. Additionally, results on a secondary dataset (reported in Appendix D) confirmed the generalizability of our methodology.
The analysis of the whole dataset allowed us to have, for the first time, a clear picture of referral motivations for this examination. This is a relevant result for public health authorities, considering the absence of structured recording of this information in the Italian NHS.
Our mapping of referrals to guidelines revealed frequent inappropriate uses, as well as numerous reasons not explicitly covered by current guidelines. While individually rare, these reasons cumulatively represent a non-negligible share of referrals, suggesting the need to expand the scope of guidelines. The ambiguity in referrals grouped under varices, where physicians often omit whether surgery is indicated, together with generic reasons like pain and vasculopathy, also points to the importance of requesting more specific information from referring physicians.
Stratification revealed significant geographical variability, suggesting that targeted interventions are needed in specific Local Health Authorities. General practitioners exhibited higher rates of inappropriate referrals compared to specialists. This might have multiple reasons, including larger patient base, more frequent encounter with less severe cases, but also a lower knowledge and/or agreement with guidelines [45, 46].
Temporal trends showed a decline in inappropriate referrals during the initial COVID-19 wave in 2020, followed by an increase in 2021 as the healthcare system normalized. This aligns with literature suggesting that during the pandemic’s peak, physicians prioritized urgent and necessary examinations to manage resource constraints effectively [47, 48, 49, 50].
Despite its strengths, this study has some limitations. Manual annotation was performed on a subset of data due to the large time required, and annotations were conducted by a single annotator, preventing inter-annotator agreement analysis. While valid for most cases, the focus on a single reason per referral overlooks referrals with multiple reasons, which could be addressed through fuzzy clustering.
Another limitation was the lack of access to patient history, which prevented a precise assessment of the appropriateness of certain referrals, such as those for varices, where the presence of a surgery often could not be verified.
It is also essential to note that this analysis does not aim to state the appropriateness of single referrals, for which a comprehensive analysis of the clinical status and history of the patient would be necessary. Instead, it aims to provide a comprehensive, automatized and accurate picture at the regional level.
In future research, we aim to expand the validation of our pipeline to other types of examinations, with larger manually annotated datasets, ensuring broader applicability across healthcare services.
6 Conclusions
This study proposed a two-step procedure for assessing the level of appropriateness of Italian referrals. We presented a case study on the venous echocolordoppler of the lower limbs in the Lombardy Region, demonstrating excellent performance on a validation set and extracting relevant insights for public health authorities. The outcomes of this analysis enabled the Lombardy Region to reinforce existing guidelines for physicians, aiming to reduce the number of inappropriate referrals. By addressing one of the key drivers of inefficiency in the healthcare system, this approach has the potential to alleviate significant burdens on the Italian NHS, lowering waiting times and enhancing the overall effectiveness of healthcare services.
[…]
List of abbreviations
  1. 1.
    NHS: National Health Service
  2. 2.
    NLP: Natural Language Processing
  3. 3.
    GP: General Practitioner
  4. 4.
    EHR: Electronic Health Record
  5. 5.
    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
  6. 6.
    GDPR: General Data Protection Regulation
  7. 7.
    IQR: Interquartile Range
  8. 8.
    ICD: International Classification of Diseases
  9. 9.
    BERT: Bidirectional Encoder Representation for Transformers
  10. 10.
    DGR: Decreto della Giunta Regionale (Regional Law)
  11. 11.
    REF: Referral
  12. 12.
    CQ: Clinical Question
Declarations
Ethics approval and Consent to participate
The need for ethics approval is deemed unnecessary according to national regulations. In particular, according to the rules from the Italian Medicines Agency (available at: https://www.gazzettaufficiale.it/eli/id/2024/08/20/24A04320/sg), retrospective studies using administrative databases do not require Ethics Committee protocol approval.
The need for consent to participate is deemed unnecessary according to national regulations. In particular, see the rules of the Italian Privacy Authority available at https://www.garanteprivacy.it/web/guest/home/docweb/-/docweb-display/docweb/9124510 and the European Union regulation n. 2016/679 (GDPR) Art. 1 c. 1 (https://gdpr.eu/article-1-subject-matter-and-objectives-overview/), Art. 4 c.1 (https://gdpr.eu/article-4-definitions/) and Recital 26 (https://gdpr-info.eu/recitals/no-26/).
Availability of data and materials
The data that support the findings of this study are available from Regione Lombardia but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available. Data are however available from the authors upon reasonable request and with permission of Regione Lombardia.
Code availability
The code is available by the authors upon request.
Competing interests
The authors declare that they have no competing interests.
Authors’ contributions
VT: investigation, formal analysis, methodology, software, data curation, writing - original draft AB: resources, data curation, conceptualization ME: resources, data curation, conceptualization OL: resources, data curation, conceptualization FI: conceptualization, methodology, formal analysis, supervision, writing - review & editing
Acknowledgments
The present research is part of the activities of “Dipartimento di Eccellenza 2023-2027”. We thank the Welfare Directorate of Regione Lombardia (Agreement between Politecnico di Milano and Regione Lombardia DGR n. XI/5301-27/09/2021) and, in particular, the working group on waiting list monitoring. We thank Dr. Paola Colombo for the consultation regarding the reasons for angiology referrals and Dr. Marco Soncini for gastroenterology referrals.
Appendix A Manual annotations
A validation set of 1000 clinical questions was manually annotated for the main disease/symptom determining the referrals. The existing guidelines from the Lombardy Region [26] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate. If none of the categories mentioned in the guidelines fit the reason for the referral, a new category representing the main disease or symptom mentioned in the referral was created. The other category was used only when it was impossible to identify a symptom or a disease determining the referral (e.g., “in-depth examination to actually establish whether there is an ongoing pathology given the symptoms present” or “gfwhdfg”).
In case of multiple diseases or symptoms, the following priority order was defined:
  1. 1.
    Reasons mentioned as appropriate in the guidelines
  2. 2.
    Reasons mentioned as inappropriate among the guidelines
  3. 3.
    Reasons not mentioned in the guidelines
If considering these priority levels, there were still multiple diseases/symptoms, the following priority order was defined:
  1. 1.
    If there is one symptom/disease which looks like the leading cause, prefer it
  2. 2.
    If there are multiple symptoms/diseases mentioned in the guidelines, choose the first one to be mentioned in the text
  3. 3.
    If there are multiple symptoms/diseases not mentioned in the guidelines, prefer those for which a category was already created, otherwise the first one to be mentioned in the text
This procedure, after four iterations, including consultation with an angiologist, defined the labels reported in Table A.1.
Table A.1: Labels derived by manual annotations on the validation set of 1000 clinical questions for venous echocolordoppler of the lower limbs. CQ = Clinical Questions, REF = Referrals.
Label # CQ # REF
Thrombosis 218 306
Venous insufficiency 200 747
Edema 104 149
(Thrombo)phlebitis 81 117
Varices 65 628
Pain 56 75
Vasculopaty 52 148
Safenectomy 27 33
Surgery 26 26
Ulcer 25 41
Paresthesis 16 25
Varices surgery 14 15
Other 13 16
Claudicatio 11 16
Swelling 9 10
Transplanctation 8 8
Teleangectasias 7 7
Cardiopathy 6 15
Dermatitis 6 12
Dimer 6 6
Ischaemia 5 5
Diabetes 4 5
Prosthesis 4 4
Cancer 3 5
Heavy Legs 3 4
Hypertension 3 6
Stent 3 2
Arthritis 2 3
Calcification 2 2
Cold legs 2 2
Dyslipidaemia 2 2
Tingling 2 3
Dialysis 1 1
Dyspnoea 1 1
Embolism 1 2
Haemorrhage 1 1
Hypercholesterolaemia 1 3
Hypertrophy 1 1
Hyposthesia 1 1
Infection 1 1
Lymphangitis 1 1
Metabolic syndrome 1 1
Multiple sclerosis 1 1
Neuropathy 1 1
Pregnancy 1 1
Smoker 1 1
Vascular encephalopathy 1 1
Appendix B Clustering
B.1 Additional details on clustering procedures
For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens. Clustering was performed with HDBscan and K-Means, obtaining the best results with K-Means. The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40.
For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7.
For HDBscan, the initial minimum size was set to 500, selected as the minimum size that led to the emergence of at least 5 clusters, and was progressively halved up to 20.
B.2 Additional clustering results
Table LABEL:tab:clusters-summ presents all the clusters identified for venous echocolordoppler of the lower limbs, with their keyword descriptions, the sizes and the labels they were assigned.
Table B.1: Summary of all clusters identified for echocolordoppler of the lower limbs, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
claudicatio; riferita claudication; referred claudication 1132 1686
claudicatio; intermittens claudicatio; intermittens claudication 653 1179
claudicatio claudicatio claudicatio 402 530
claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudicatio 60 126
claudicatio; arti; inferiori claudicatio; limbs; lower claudicatio 41 49
dermatite dermatitis dermatitis 421 474
diabete diabetes diabetes 1540 3249
diabete; mellito diabetes; mellitus diabetes 534 1847
diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
diabetico; diabetica diabetes; diabetic diabetes 209 165
dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
dialisi dialysis dialysis 110 222
dimero dimer dimer 1343 1601
dispnea dyspnoea dyspnoea 104 171
dolore pain pain 6343 12051
deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
polpaccio; dolore calf; pain pain 707 1753
algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
dolore; popliteo pain; popliteal pain 35 117
algia; covid; gamba pain; covid; leg pain 30 63
dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
crampi; notturni cramps; nocturnal pain 6 5
edema edema edema 11686 20723
edemi; declivi edema; declivities edema 9491 16473
edema; arti; edemi edema; limbs; edemas edema 3712 7867
arti; inferiori; edemi limbs; lower; edemas edema 528 2486
edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
edema; arti; inferiori edema; limbs; lower edema 899 2350
edema; perimalleolare edema; perimalleolar edema 1039 2214
edema; natura; determinare edema; nature; determine edema 1027 2199
edemi; perimalleolari edema; perimalleolar edema 1190 2151
edema; malleolare edema; malleolar edema 935 1136
edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
edema; flebite edema; phlebitis edema 217 297
arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
arti; inferiori; edema limbs; lower; edema edema 80 168
arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
edema; inferiori; arti edema; lower; limbs edema 74 163
edema; sinistra edema; left edema 61 108
claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
linfedema; arti lymphedema; limbs edema 23 24
emorragia haemorrhage haemorrhage 18 29
formicolio tingling tingling 43 170
freddo cold cold 49 147
fumatore smoker smoker 121 193
gravidanza pregnancy pregnancy 92 225
insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
insufficienza; venosa venous; insufficiency venous insufficiency 0 0
arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
venosa; cronica venous; chronic venous insufficiency 0 0
insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
ipoestesia hypoesthesia hypoesthesia 31 80
ischemia ischaemia ischaemia 517 617
parestesie paresthesias paresthesias 1834 3965
parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
pesantezza heaviness heaviness 215 475
arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
safenectomia saphenectomy saphenectomy 1988 4177
safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
sindrome metabolica metabolic syndrome metabolic syndrome 12 48
stent stent stent 301 612
tromboflebite thrombophlebitis thrombophlebitis 7073 17415
flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
trombosi thrombosis thrombosis 17060 43558
trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
trombosi; venosa venous; thrombosis thrombosis 1325 3337
venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
poplitea; trombosi popliteal; thrombosis thrombosis 607 642
venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
safena; trombosi saphenous vein; thrombosis thrombosis 301 263
post; trombotica post; thrombotic thrombosis 36 73
tumefazione swelling swelling 533 741
tumefazione; gamba swelling; leg swelling 172 448
destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
ulcera; vascolare ulcer; vascular ulcer 828 3861
ulcera ulcer ulcer 1258 2411
varici varices varices 13999 35939
arti; inferiori; varici limbs; lower; varices varices 7018 18157
venose; varici venous; varices varices 4957 13554
varicosita; stasi varicosities; stasis varices 957 2701
varici; arti; inferiori varices; limbs; lower limbs varices 219 277
vene; varicose; algie veins; varicose; algie varices 121 194
safena; varici saphenous veins; varicose varices 47 136
varicosita; varicosita’ varicosities; varicositis varices 86 97
vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
varici; sospette varicose veins; suspected varices 34 63
varicosita; gamba varicosities; leg varices 80 63
varici; intervento varicosities; surgery varices intervention 3144 7126
vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
vasculopatia vasculopathy; peripheral vasculopathy 769 1886
vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
arteriopatia arteriopathy vasculopathy 20 25
natura; determinare; periferica nature; determine; peripheral other 3978 10039
natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
arti; inferiori limbs; lower other 2190 5165
safena; incontinenza saphenous; incontinence other 288 782
ipertrofia hypertrophy other 506 591
tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
embolia; polmonare embolism; pulmonary other 110 174
femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
stenosi; venoso stenosis; venous other 30 49
dislipidemia; ipertensione dyslipidaemia; hypertension other 25 44
aorta; sottorenale aorta; subrenal other 25 29
Table B.1: *
Appendix C Pre-processing results
Table C.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table C.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing.
Table C.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
Abbr Freq Expansion Expansion [ENG]
dx 41942 destra right
sx 35400 sinistra left
tvp 35392 trombosi venosa profonda deep vein thrombosis
inf 26644 inferiore inferior
sin 8490 sinistra left
pz 6100 paziente patient
ndd 4555 di natura da determinare of a nature to be determined
$a$ 4079 arti limbs
tvs 3650 trombosi venosa superficiale superficial venous thrombosis
sn 3055 sinistra left
x 2210 per for
k 2170 cancro cancer
ivc 2124 insufficienza venosa cronica chronic venous insufficiency
$i$ 1768 inferiori inferior
tep 1695 tromboembolia polmonare pulmonary thromboembolism
nao 1291 nuovi anticoagulanti orali new oral anticoagulants
ca 1262 cancro cancer
ds 1001 destra right
paz 959 paziente patient
tao 924 terapia anticoagulante orale oral anticoagulant therapy
vgs 911 vena grande safena great saphenous vein
tv 875 trombosi venosa venous thrombosis
tp 852 trombosi profonda deep thrombosis
s 819 sindrome syndrome
dm 800 diabete mellito diabetes mellitus
bil 792 bilaterale bilateral
irc 650 insufficienza renale cronica chronic renal failure
ins 620 insufficienza failure
sup 608 superiore upper
fa 592 fibrillazione atriale atrial fibrillation
v 576 venosa venous
iv 497 insufficienza venosa venous insufficiency
spt 469 sindrome post trombotica post-thrombotic syndrome
npl 451 neoplasia neoplasm
iii 444 inferiori inferior
ep 421 episodio episode
tx 364 trapianto transplantation
ctr 351 controllo control
art 327 tromboembolismo venoso venous thromboembolism
tpv 326 trombosi profonda venosa deep vein thrombosis
vs 307 vena safena saphenous vein
cr 297 cronica chronic
$g$ 269 giorni days
tev 256 tromboembolia venosa venous thromboembolism
pta 251 angioplastica percutanea transluminale percutaneous transluminal angioplasty
ven 235 venosa venous
int 227 intervento surgery
mvc 225 malattia venosa cronica chronic venous disease
ev 220 eventuale possible
$t$ 203 tibio-tarsica tibio-tarsal
Table C.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.
Original Word Corrected Word Corrected Word [ENG]
iprtensioone ipertensione hypertension
stwnosi stenosi stenosis
areomasica ateromasia atheromasia
aduta caduta fall
clacifiche calcifiche calcifications
lipitimiche lipotimie lipothymias
esignificative significative significant
veriginosa vertiginosa vertiginous
lipotimoco lipotimici lipotimics
emicrani emicrania migraine
cardiopatiia cardiopatia heart disease
30anni anni years
sinndrome sindrome syndrome
vertigicnosa vertiginosa vertiginous
eriosa arteriosa arterial
srteriopatia arteriopatia arteriopathy
carotgidee carotidea carotid
nvertiginosa vertiginosa vertiginous
nlipotimia lipotimici lipothymic
pregressotia pregresso previous
priomo primo first
dislipemua dislipidemia dyslipidaemia
caroltidea carotidea carotid
iportesa ipertesa hypertensive
cerebricolare cerebrovascolare cerebrovascular
neurologoin neurologo neurologist
cronarico coronarica coronary
arteiopatia arteriopatia arteriopathy
bilaterati bilaterale bilateral
aritimica aritmica arrhythmic
bpcoe bpco bpco
extrasisoliche extrasistolia extrasystole
paiente paziente patient
iprtnsione ipertensione hypertension
teriopati arteriopatia arteriopathy
caprtidea carotidea carotid
soszpetta sospetta suspected
dislidemica dislipidemia dyslipidaemia
lipotimicio lipotimici lipothymics
epeisodi episodi episodes
ipertrigliceridiemia ipertrigliceridemia hypertriglyceridaemia
ertiginosa vertiginosa vertigo
iperpensione ipertensione hypertension
neopla neoplasia neoplasia
oderata moderata moderate
cardiopataia cardiopatia cardiopathy
polidistrettuae polidistrettuale polydistrict
dislipidemiaecocolo dislipidemico dyslipidaemic
ipwrtensone ipertensione hypertension
acugene acufene tinnitus
Appendix D Additional results on gastroenterology exam
Additional results related to the flexible endoscope colonoscopy with possible biopsies examination are reported in this section. The dataset for this exam consisted of all referrals filled in in the Lombardy Region between 2019 and 2021, for a total of 407,949 referrals. A random sample of 1,000 distinct clinical questions (corresponding to 3,556 referrals) was annotated in the same way as for the case study presented in the main manuscript and used to validate the results.
Considering 2015 guidelines from Regione Lombardia for this type of exam [27], we marked colon cancer familiarity/prevention and chronic inflammatory bowel disease as likely appropriate. Guidelines state that, in these cases, physicians should carefully consider the patient history and condition to establish if and when the exam should be performed.
Table D.1 reports the detailed results per category, while Table D.2 the summary metrics considering the categories and the appropriate/inappropriate levels. Table D.3 presents the final results on the full dataset, which are stratified by year, type of prescribing physician and LHU in Figure D.1.
Table D.4 details the labels from manual annotation on the test set.
Table D.1: Detailed results of the cluster mapping for each category, on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF). P=Precision, R=Recall, F1=F1-Score.
Category P-CQ [%] R-CQ [%] F1-CQ [%] # CQ P-REF [%] R-REF [%] F1-REF[%] # REF
Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
Family history / prevention 97.56 86.33 91.60 139 98.57 91.39 94.85 302
Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
Chronic inflammatory bowel disease 100.00 89.29 94.34 56 100.00 88.37 93.83 86
Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
Table D.2: Summary metrics on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF), considering annotations categories or only appropriate/inappropriate/likely appropriate/undetermined. P=Precision, R=Recall, F1=F1-Score.
P-CQ [%] R-CQ [%] F1-CQ [%] P-REF [%] R-REF [%] F1-REF [%]
Annotation categories 90.42 88.78 88.60 93.59 92.70 92.67
Appropriateness levels 92.05 90.96 90.88 94.66 93.96 94.05
Table D.3: Summary of referrals’ clusters for flexible endoscope colonscopy, mapped to all labels defined in manual annotations. Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
Cluster Label # REF %
polyps 81,111 19.88
occult blood 76,487 18.75
colon cancer familiarity / prevention 43,615 10.69
abdominalgia 31,316 7.68
colon cancer 30,434 7.46
anaemia 23,133 5.67
chronic inflammatory bowel disease 21,470 5.26
constipation 15,109 3.70
diverticulosis 15,058 3.69
chronic diarrhoea 13,533 3.32
altered alvo 10,933 2.68
haemorrhoidal disease 5,379 1.32
diverticulitis 4,669 1.14
suspected colic lesion detected by other investigation technique 3,671 0.90
irritable bowel syndrome 2,023 0.50
surgery 797 0.20
lynch syndrome 787 0.19
other 28,424 6.98
Figure D.1: Prescription appropriateness for flexible endoscope colonscopy grouped by LHA (A), type of prescribing physicians (B) and year (C)
Table D.4: Labels derived by manual annotations on the validation set of 1000 clinical questions for flexible endoscope colonoscopy.
Label # CQ # REF
polyps 158 344
bleeding 122 954
colon cancer 107 187
abdominalgia 104 214
colon cancer familiarity 82 141
anaemia 65 115
prevention 57 161
occult blood 36 789
constipation 36 139
diverticulosis 33 48
altered bowel 30 65
colitis 28 40
chronic diarrhoea 28 107
diverticulitis 14 28
crohn’s disease 12 16
haemorrhoidal disease 9 30
chronic inflammatory bowel disease 8 15
ulcerative rectocolitis 8 15
surgery 5 6
intestinal obstruction 5 10
proctorrhagia 5 5
dyspepsia 4 5
weight loss 3 8
irritable bowel syndrome 3 6
lymphoma 3 4
transplant 3 15
other 2 2
colonpathy 2 7
enteritis 2 3
proctitis 2 4
rectal prolapse 2 12
lynch syndrome 2 2
angiodysplasia 1 1
appendicitis 1 1
celiac disease 1 1
intestinal loops distension 1 1
hepatopathy 1 1
protruding formation 1 1
gastritis 1 2
ileitis 1 1
hypercholesterolaemia 1 1
splenic flexure thickening 1 3
rectal wall thickening 1 2
sigma wall thickening 1 1
behcet’s disease 1 1
liver disease 1 1
pancreatitis 1 1
proctalgia 1 3
sigmoiditis 1 1
suspected lesion 1 33
splenic fissure stenosis 1 1
rectal tenesmus 1 2

### v2

@@
 Abstract
-Objective: This study proposes a Natural Language Processing pipeline to evaluate prescription appropriateness in Italian referrals, where reasons for prescriptions are recorded only as free text, complicating automated comparisons with guidelines.
-The pipeline aims to derive, for the first time, a comprehensive summary of the reasons behind these referrals and a quantification of their appropriateness.
-While demonstrated in a specific case study, the approach is designed to generalize to other types of examinations.
-Methods: Leveraging embeddings from a transformer-based model, the proposed approach clusters referral texts, maps clusters to labels, and aligns these labels with existing guidelines.
-We present a case study on a dataset of 496,971 referrals, consisting of all referrals for venous echocolordopplers of the lower limbs between 2019 and 2021 in the Lombardy Region.
-A sample of 1,000 referrals was manually annotated to validate the results.
-Results: The pipeline exhibited high performance for referrals’ reasons (Prec=92.43%, Rec=83.28%) and excellent results for referrals’ appropriateness (Prec=93.58%, Rec=91.52%) on the annotated subset.
-Analysis of the entire dataset identified clusters matching guideline-defined reasons - both appropriate and inappropriate - as well as clusters not addressed in the guidelines.
-Overall, 34.32% of referrals were marked as appropriate, 34.07% inappropriate, 14.37% likely inappropriate, and 17.24% could not be mapped to guidelines.
-Conclusions: The proposed pipeline effectively assessed prescription appropriateness across a large dataset, serving as a valuable tool for health authorities.
-Findings have informed the Lombardy Region’s efforts to strengthen recommendations and reduce the burden of inappropriate referrals.
+Background: Assessing the appropriateness of diagnostic referrals is critical for improving healthcare efficiency and reducing unnecessary procedures.
+However, this task becomes challenging when referral reasons are recorded only as free text rather than structured codes.
+In the Italian National Health Service (NHS), referral motivations are systematically entered as unstructured text, making large-scale evaluation against official guidelines difficult to automate and scale.
+To address this gap, we propose a fully unsupervised Natural Language Processing (NLP) pipeline capable of extracting and evaluating referral reasons without relying on labelled datasets, which are costly and time-consuming to produce.
+Methods: Our pipeline leverages Transformer-based embeddings pre-trained on Italian medical texts to semantically cluster referral reasons and assess their alignment with appropriateness guidelines.
+It operates in an unsupervised setting and is designed to generalize across different examination types.
+We analyzed two complete regional datasets from the Lombardy Region (Italy), covering all referrals between 2019 and 2021 for venous echocolordoppler of the lower limbs (ECD; n=496,971) and flexible endoscope colonoscopy (FEC; n=407,949).
+The pipeline was developed on the ECD dataset and tested on the FEC dataset.
+For both, a random sample of 1,000 referrals was manually annotated to measure performance.
+Results were also stratified by year, region, and physician type.
+Results: The pipeline achieved high performance in identifying referral reasons (Precision=92.43% (ECD), 93.59% (FEC); Recall=83.28% (ECD), 92.70% (FEC)) and appropriateness (Precision=93.58% (ECD), 94.66% (FEC); Recall=91.52% (ECD), 93.96% (FEC)).
+Inappropriate referrals were significantly more prevalent for ECD (34.1%) than for FEC (5.3%).
+The analysis revealed geographical, temporal, and physician type variations, including a drop in inappropriate referrals during the COVID-19 pandemic in 2020.
+Additionally, several referral motivations not covered by current guidelines were identified, highlighting areas for policy improvement.
+These findings contributed to updates to the regional guidelines in Lombardy.
+Conclusions: This study presents a robust, scalable, unsupervised NLP pipeline for assessing referral appropriateness from unstructured texts in large, real-world datasets.
+It demonstrates how such data can be effectively leveraged, providing public health authorities with a deployable AI tool to monitor practices and support evidence-based policy.
 keywords:
-Natural language processing, Referrals, Prescription appropriateness, Public healthcare system, Quality of healthcare
-1 Background and significance
-In Italy, the National Health System (NHS) is the primary provider of healthcare services, offering universal coverage, with decentralized administration at the regional level [1].
-Despite a strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, leading many citizens to turn to private providers [2].
-One significant contributor to these delays is the issue of prescription appropriateness, a concern shared by many advanced healthcare systems [3, 4].
-Inappropriate prescriptions have a threefold negative impact on the healthcare system: waste of resources, increased waiting times, and reduced effectiveness of the care provided to citizens, who face delays in obtaining accurate diagnoses [5].
-In the Italian NHS, referrals from general practitioners (GPs) or specialized physicians are necessary to book examinations.
-Assessing prescription appropriateness requires understanding the reasons underlying these referrals.
-When Italian physicians fill out a referral, they have to use a standardized form, providing the reason for the prescription in a specific free-text field called Clinical Question.
-The unstructured nature of this field, coupled with the lack of standardized alternatives, makes it challenging to analyze.
+Natural language processing, Artificial Intelligence, Machine Learning, Referrals, Unsupervised learning, Examination appropriateness, Guideline adherence, Public healthcare system, Quality of healthcare
+pacs:
+MSC Classification]68T50
+1 Background
+In Italy, the National Health Service (NHS) is the primary provider of healthcare services, offering universal coverage, with administration decentralized at the regional level [1].
+Despite its strong performance in international rankings, the Italian NHS has faced growing challenges in providing timely access to healthcare services in recent years, with increasing numbers of citizens turning to private providers [2].
+One significant contributor to these delays is the issue of referrals appropriateness, a concern shared by many advanced healthcare systems [3, 4].
+Inappropriate referrals lead to resource waste, increased waiting times, and delays in accurate diagnoses, ultimately reducing the effectiveness of the healthcare system [5].
+Within the Italian NHS, referrals from general practitioners (GPs) and specialists are required to book diagnostic examinations, making it crucial to assess the reasons behind these referrals to ensure their appropriateness.
+When issuing a referral, Italian physicians must complete a standardized form, including a free-text field known as the Clinical Question, where they specify the reason for the prescription.
+The unstructured nature of this field, coupled with the absence of standardized alternatives, poses a major challenge for systematic analysis.
 Addressing this issue requires Natural Language Processing (NLP) techniques.
-The number of studies analyzing texts of Italian referrals is very limited.
+In this study, we frame the task as a clustering problem: clinical questions are grouped using Transformer-based embeddings [6] to capture semantic similarity, and clusters are subsequently mapped to guideline-derived categories of appropriateness.
+Despite its importance, research on NLP-driven analysis of Italian referrals is limited.
 Venturelli et al.
-[6] used proprietary commercial software (Clinika VAP [7]), for which the NLP models or algorithms are not disclosed.
-This software provides an appropriate/inappropriate label for each referral, enabling the quantification of inappropriate referrals.
-However, it does not provide a clustered analysis of the referrals, with a quantification of the different prescription reasons.
-To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [8], which focused on extracting and analyzing follow-up waiting times from clinical questions to investigate timely access to care.
-While several studies have evaluated prescription appropriateness for medications within the Italian NHS using structured data or manual text analysis (e.g., [9, 10, 11]), research on the appropriateness of examinations is more limited [12, 13] and has not incorporated natural language processing.
-A few studies developed decision-support systems for guideline adherence, both in primary care [14, 15] and hospital settings [16, 17].
-These tools assist physicians through closed-form questions, possibly complemented by keyword matching on the text they write.
+[7] used proprietary commercial software (Clinika VAP [8]) to classify referrals as appropriate or inappropriate, but without explicitly identifying the reasons for prescriptions.
+To the best of our knowledge, the only other study applying NLP on Italian referrals’ texts is [9], which focused on extracting and analyzing follow-up waiting times from clinical questions rather than assessing referrals appropriateness.
+Most previous studies on appropriateness in the Italian NHS relied on structured data or manual text analysis, focusing on medication prescriptions rather than medical examinations [10, 11, 12].
+Research on the appropriateness of examinations is more limited [13, 14] and has not incorporated textual data.
+Other studies focused on the development of decision-support systems for guideline adherence, both in primary care [15, 16] and hospital settings [17, 18].
 While effective in specific scenarios [19, 20], such tools are not suited for retrospective analyses of their performance or for quantifying the prevalence of appropriate or inappropriate referrals.
-Beyond the Italian context, Haddad et al.
-[20] analyzed Russian referrals’ texts for prescription appropriateness using labelled data; Sagheb et al.
+Internationally, several studies have employed NLP and rule-based systems to assess referral appropriateness, often in disease-specific contexts.
+Haddad et al.
+[21] analyzed Russian referrals’ texts for appropriateness using labelled data; Sagheb et al.
 [22] designed a rule-based algorithm to verify guidelines adherence for asthma patients from EHR documents; similarly in [23] a rule-based approach was applied to assess guideline adherence for gout patients.
@@
 Villeana et al.
-[25], analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
-These studies, however, are often disease-specific or reliant on labelled data or proprietary software, limiting their broader applicability.
-2 Objective
-This study proposes an NLP-based pipeline to assess the appropriateness of Italian referrals, developing a model designed for unsupervised analysis across different types of medical examinations.
-In 2015, the Lombardy Region, the most populated Italian region with approximately 10 million inhabitants - akin to a small state - issued two resolutions (DGRs) focusing on prescription appropriateness in angiology [26] and gastroenterology [27], two fields where the literature reports high rates of inappropriate prescriptions, both in Italy and internationally [28, 29, 30, 31, 32].
-These resolutions established guidelines specifying appropriate and inappropriate clinical questions for a variety of examinations.
-Building on these efforts, we analyzed examinations covered by these resolutions, contributing to the issuance of a new resolution in December 2023 that reinforced the previous guidelines [33].
-This paper focuses on a case study of one of these examinations: the venous echocolordoppler of the lower limbs.
-In particular, we analyzed a dataset consisting of all referrals prescribed in the Lombardy Region for this exam between 2019 and 2021, for a total of 496,971 referrals, encompassing 229,509 unique clinical questions.
-The analysis presented in this study aims to verify the effectiveness of the 2015 resolutions in enforcing prescriptive appropriateness, providing for the first time an overview of the main reasons why these examinations are prescribed, with an approximate quantification of the number of referrals for each disease/symptom and their alignment on existing guidelines, where possible.
-Given the absence of labelled data and the need to uncover reasons not addressed in guidelines, we framed this as a clustering problem.
-In particular, our pipeline clusters clinical questions by exploiting the embeddings produced by a fine-tuned BERT model [34].
-These clusters are summarized and, where possible, mapped to guidelines to determine whether they are appropriate, inappropriate, or not covered by the guidelines.
-A subset of 1,000 distinct clinical questions was manually annotated to validate the pipeline.
-Results are reported for this validation subset to verify the correctness of the pipeline output and for the full dataset to analyze prescription appropriateness at the regional level, including stratifications by year (pre-, during, and post-COVID pandemic), geographical area, and type of prescribing physician.
-Additional results for another type of examination in the gastroenterological domain are included in Appendix D.
-The main contributions of this study are the following:
-  1. 1.
-    The first open-source, fully unsupervised pipeline for assessing prescription appropriateness from Italian referrals, applicable to any type of examination.
-  2. 2.
-    The first region-wide analysis in Italy of referrals’ reasons and their appropriateness.
-3 Material and Methods
+[26] analyzed Chilean referrals’ texts using labelled data to binary classify them for correct priority classes.
+However, all these approaches rely on labeled data or proprietary software, limiting their adaptability and scalability across different examination types.
+To address these gaps, this study proposes the first fully unsupervised NLP pipeline for analyzing appropriateness in Italian referrals.
+The pipeline is applicable to any examination for which guidelines define appropriate and inappropriate referral reasons.
+We developed this approach on a first case study related to venous echocolordoppler of the lower limbs (ECD) and we present results also on a second case study related to flexible endoscope colonoscopy (FEC).
+Both datasets include all referrals issued during three years in the entire the Lombardy Region, the most populated Italian region, with approximately 10 million inhabitants - akin to a small state.
+These two examinations were selected because both are characterized by a high prevalence of inappropriate referrals [27, 28, 29, 30] and had already been targeted by Lombardy’s 2015 resolutions on angiology and gastroenterology [31, 32].
+However, no comprehensive study has assessed the impact of those resolutions at scale.
+Our analysis provides, for the first time, a regional-scale evaluation of referral appropriateness and of guideline effectiveness, which informed the Lombardy Region’s December 2023 resolution [33].
+This work thus provides both a methodological contribution, by introducing a scalable NLP pipeline for appropriateness assessment, and an applied contribution, by generating evidence that informed regional policies.
+The remainder of this paper is structured as follows: Section 2 describes the datasets and the pipeline, Section 3 presents the results, Section 4 addresses implications, limitations, and future developments; and Section 5 concludes.
+2 Methods
 2.1 Data
-3.1.1 Referrals dataset
-The dataset analyzed in this study is extracted from all referrals for venous echocolordoppler of the lower limbs prescribed in the Lombardy Region between 2019 and 2021 (496,971 referrals).
+2.1.1 Referrals datasets
+The datasets analyzed in this study are extracted from all referrals issued in the Lombardy Region between 2019 and 2021 for ECD and FEC, totalling 496,971 referrals for ECD and 407,949 for FEC.
 Each referral includes the clinical question, the year, the Local Health Authority (LHA) of the prescribing physician and the type of prescribing physician.
-Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs represent the eight healthcare administrative areas within the Lombardy Region.
+Physician types are categorized as general practitioners (GPs) or specialized physicians, while LHAs are the eight healthcare administrative areas within the Lombardy Region.
 Figure 1 illustrates the distribution of referrals among these groups.
-The analysis focuses on the clinical question, a free-text field typically containing one or few sentences.
+Figure 1: Referrals distributions across years (A, B), type of prescribing physicians (C, D) and LHA (E, F) for ECD (A,C,E) and FEC (B,D,F).
+The analysis focuses on the clinical question (CQ), a free-text field typically containing one or a few sentences.
 These texts exhibit frequent spelling and typing errors, heavy use of abbreviations, inconsistent punctuation, and non-standard capitalization.
-The dataset includes 229,509 distinct clinical questions, but only 15% appear more than once.
-The median length is 71 characters (IQR=39).
-Only 189 (0.08%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
-The clinical questions should not contain personal information about patients or physicians.
-A preliminary check was performed by the Lombardy Region IT provider, masking any suspect personal information before data extraction.
+The dataset for ECD includes 229,509 distinct CQ, with only 15% appearing more than once, while the dataset for FEC includes 89,734 distinct CQ, with 41% appearing more than once.
+The median length for ECD is 71 characters (IQR=39), while for FEC it is 57 (IQR=20).
+Only 189 for ECD (0.08%) and 137 for FEC (0.15%) contain a meaningful ICD code, underscoring the necessity of employing NLP for this analysis.
+The clinical questions do not contain personal information about patients or physicians.
 2.1.2 Manual annotations
-To validate the identified clusters, a randomly selected subset of 1,000 distinct clinical questions was manually labelled.
+To validate the pipeline, a randomly selected subset of 1,000 distinct clinical questions was manually labelled for both ECD and FEC.
 The labels correspond to categories representing diseases or symptoms that prompted the referral.
-During an initial review of this subset, ambiguous cases were identified, and an angiologist was consulted to solve doubts and align clusters with prescription guidelines.
+During an initial review of this subset, ambiguous cases were identified, and an angiologist and a gastroenterologist were consulted to resolve these doubts and align the clusters with prescription guidelines.
 Annotation guidelines were subsequently refined, and the entire subset was annotated accordingly.
-Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled dataset was error-free.
-More details on the annotations are provided in Appendix A.
+Two other reviews of the annotations corrected remaining errors and clarified uncertainties, ensuring the final labelled datasets were error-free.
+More details on the manual annotations are provided in Additional File 1, Section A.
 2.1.3 Prescription guidelines
-The guidelines used as a reference to assess referral appropriateness are extracted from [26].
-In particular, we extracted the appropriate and inappropriate diseases and symptoms for the venous echocolordoppler of the lower limbs.
-An English translation is reported in Figure 2.
-Figure 1: Referrals distributions across years (A), type of prescribing physicians (B) and LHA (C)
-Figure 2: Guidelines for referrals for venous echocolordoppler of the lower limbs in the Lombardy Region. Translated from the Appendix 3 of [26].
-3.2 Pipeline for prescription appropriateness
+The guidelines used as a reference to assess referral appropriateness are extracted from [31] for ECD and from [32] for FEC.
+In particular, we extracted the appropriate and inappropriate diseases and symptoms, considering separately also cases that are more likely to be appropriate or more likely to be inappropriate, depending on details that might not be available in clinical questions.
+English schematized summaries of these guidelines are reported in Figure 2.
+Figure 2: Guidelines for referrals for ECD and FEC in the Lombardy Region. Derived from the the Appendix 3 of [31] and the Appendix 2 of [32], translated into English.
+2.2 Evaluation of prescription appropriateness
 In this section, we present our pipeline to evaluate prescription appropriateness. The process can be divided into two main steps:
@@
 Because of this, the problem was addressed as a clustering problem.
-Figure 3 depicts blocks composing our pipeline, which are detailed below.
+Figure 3 depicts our pipeline with the blocks that compose it, and below, we detail each block.
 Figure 3: Schema of the pipeline for assessing referrals reasons and appropriateness
@@
 The use of uppercase letters does not have any particular meaning, and the use of abbreviations is very frequent.
-Even in the single type of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms.
-In the manually labelled subset, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
-To address these issues, suitable pre-processing steps were applied:
+Even in the two types of examination analyzed in this study, multiple abbreviations can indicate the same disease/symptoms.
+In the manually labelled subsets, 15% of clinical questions were marked as having typing errors and 68% as containing abbreviations.
+To address these issues, suitable pre-processing steps are applied:
   1. 1.
@@
     (partial) typos correction
-We removed all punctuation and lowercase sentences due to the inconsistent use of these elements.
-[…]
-[…]
-Examples of pre-processing effects are reported in Appendix C.
+We remove all punctuation and lowercase sentences due to the inconsistent use of these elements.
+For abbreviation expansion, we extract all the unique words with length $\leq~3$, and we identify those corresponding to abbreviations of medical terms, defining a vocabulary for their replacement in the texts.
+To fix the typing mistakes, all words appearing with frequency $1\%$ and having another word similar to them with frequency $1\%$ are selected.
+These words are considered misspelt, and the most similar word with frequency $1\%$ is chosen as their correction.
+The similarity is measured with edit distance, and the minimum threshold is chosen by minimizing errors in a manually reviewed sample of 100 corrections.
+Examples of pre-processing effects and an ablation analysis are reported in Additional File 1, Section B.
 2.2.2 Text embedding
 To appropriately cluster the clinical questions, it is necessary to produce a representation capturing the central semantics of these texts without getting confused by their widely present noise.
-Because of this, vocabulary-based representations, such as TF-IDF [35], cannot be expected to work well for this task.
-The main alternatives are static and contextual neural network-based embeddings.
+Considering the many synonyms and composite terms of the medical jargon used in these texts, vocabulary-based representations, such as TF-IDF [34], cannot be expected to work well for this task.
 Static word embeddings, such as Word2Vec [35], can capture the semantic similarities between words with similar meanings (e.g., vasculopathy - arteriopathy, phlebitis - thrombophlebitis) differently from vocabulary-based representations.
 Nevertheless, they might have more difficulty distinguishing between relevant and irrelevant tokens in our sentences since they do not consider the context in which tokens appear.
-On the contrary, context-based embeddings produced by transformer models consider the surrounding context while embedding each token, and have often shown superior performances in the recent clinical NLP literature [37, 38, 39, 40].
+On the contrary, context-based embeddings produced by transformers model consider the surrounding context while embedding each token and have often shown superior performances in the recent clinical NLP literature [36, 37, 38, 39].
 The problem is particularly challenging in our case due to the limited number of embedding models available for the Italian language, especially for the clinical domain.
-To this purpose, the same model used in [41] was adopted to cluster sentences extracted from Italian referrals, and we verified that even with clinical questions, its performances are superior to other models.
+To this purpose, we adopt the same model developed in [40] to cluster sentences extracted from Italian discharge summaries, and we verify that even with referrals’ clinical questions, its performances are superior to other models.
 This model consists of a fine-tuned version of the last Italian version of BERT, Umberto [41], on a dataset of publicly available medical documents in Italian and other languages (translated into Italian) [42].
 2.2.3 Clustering
-HDBscan was adopted to cluster the embeddings since it does not require eliciting in advance the number of clusters and it allows a not complete assignment of items to clusters.
-HDBscan requires defining the minimum size of the clusters to identify.
-We verified that the best results are achieved when this parameter is progressively reduced while iterating the clustering on the previously unclustered data.
-This approach reduced the number of unclustered points, addressing the varying cluster sizes in our dataset.
-An additional issue in the clustering procedure was related to the size of our dataset.
-On a machine with 32 GB of RAM, it was not possible to run the clustering algorithm on the whole dataset.
-Because of this, we had to divide it into chunks of 30,000 clinical questions and cluster them separately.
-The subsequent summarization and mapping steps merged the clusters found in different chunks and corresponding to the same diseases/symptoms.
+Some studies have shown that HDBSCAN often outperforms other clustering algorithms—such as K-Means, hierarchical clustering, and DBSCAN-when applied to text embeddings, particularly when preceded by a dimensionality reduction step [43, 44].
+Text embeddings typically lie on complex, non-linear manifolds and may form clusters with varying densities and irregular shapes.
+Algorithms like K-Means assume convex cluster geometries and require predefining the number of clusters, which is impractical for our task.
+Hierarchical clustering is sensitive to noise and outliers and lacks the ability to leave out unclustered data points, often producing forced or less meaningful groupings.
+In contrast, HDBSCAN does not require the number of clusters to be specified in advance and allows some data points to remain unclustered, making it a better fit for our scenario.
+Based on these advantages and on a comparison of the results on the ECD dataset, we selected HDBSCAN for the clustering step.
+HDBSCAN requires specifying some hyperparameters, including the minimum cluster size, the minimum number of samples in the neighborhood of a core point (min_samplesmin\_samples), and the epsilon.
+To better accommodate the varying sizes of clusters in our dataset, we adopted a recursive strategy: the minimum cluster size was iteratively reduced, halving at each step, and applied only to data points that remained unclustered in the previous iteration.
+For the dimensionality reduction, we compared PCA, UMAP and no dimensionality reduction.
+We performed a grid search and we selected the configuration with PCA with 80% of variance threshold, reducing the minimum cluster size from 500 to 10, with min_samples=5min\_samples=5 and $\epsilon=0.1$.
+The complete grid search results, together with comparisons to k-means and agglomerative hierarchical clustering, are reported in Additional File 1, Section C.
+Due to resource constraints, it was not feasible to run HDBSCAN on the entire datasets.
+Therefore, we divided the data into chunks of 30,000 clinical questions and clustered them independently.
+In later stages, the summarization and mapping procedures allow to merge clusters from different chunks that correspond to the same diseases or symptoms.
 2.2.4 Cluster summarization
@@
 To automatize this step, we extracted keywords from each cluster.
-The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold until at least two keywords were found.
+The most frequent words in each cluster were initially selected, considering a decreasing frequency threshold from 80% until at least two keywords were found.
 Among these, we excluded those with high frequency in at least half of the clusters.
-This allowed us to exclude not only stop words but also frequent words for the specific exam (e.g., leg, lower limbs, right, left in our case study).
+This allowed us to exclude not only stop words but also frequent words for the specific examination we are analyzing.
 2.2.5 Mapping clusters to guidelines
-After summarization, the keyword representation of the clusters was used to map them to the guidelines. The process involved:
+After summarization, the keyword representation of the clusters is used to map them to the guidelines. The process involves:
   1. 1.
@@
     Comparing cluster embeddings with guideline embeddings using cosine similarity.
-Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50absent0.50>0.50> 0.50.
+Clusters were assigned to the guideline entry with the highest similarity, provided there is one with similarity >0.50>0.50.
 2.3 Experiments and Evaluation
-3.3.1 Pipeline evaluation
-To validate the pipeline, we compared its output with labels from the manually annotated subset.
-The pipeline was executed on the entire dataset, and results for the manually labelled portion were extracted.
-This approach ensures fairness, as running the pipeline solely on the labelled subset (1,000 records) would limit its ability to form meaningful clusters due to the restricted data size.
-Manual annotations identified 49 distinct labels, with only 13 occurring more than 10 times.
-For evaluation purposes, these 13 labels were retained for comparison, while the remaining labels were grouped into the other category, which we expect to align with unclustered elements.
-The results were evaluated in terms of precision, recall, and f1-score.
-Metrics were calculated at two levels: at aggregate level to evaluate the overall pipeline performance across all labels, and at per-label level to evaluate performance for individual labels.
-Additionally, we weighted clinical questions by their corresponding referral counts, incorporating the dataset’s real-world distribution into the analysis.
-A second type of evaluation is performed to evaluate the quality of prescription appropriateness labels.
-In this case, we grouped the 13 labels into four categories: appropriate, inappropriate, likely inappropriate and undetermined.
-In particular, we marked as likely inappropriate some clusters for which the indication was not completely clear (see Results section for details on this), and as undetermined those that were not covered by the guidelines.
-In both cases, we compared our pipeline with other types of text representations/clustering algorithms.
-We considered clustering results obtained by embeddings from the original Umberto model, by embeddings from an Italian Word2Vec model [44], by TF-IDF representations and by the LDA algorithm.
-See Appendix B for details.
-3.3.2 Analysis of the entire dataset
-The pipeline is applied to the entire dataset, with results that provide a complete quantification of referrals for each identified category, together with their appropriateness.
-Moreover, we present stratified results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs specialists).
+In this section, we detail the experimental settings and the evaluation procedure for our pipeline, which are summarized in Figure 4.
+Figure 4: Summary of the experiments and evaluation procedure
+2.3.1 Experimental Setup and Constraints
+All experiments were conducted in an offline environment with limited computational resources: 16 GB of RAM, 16 cores Intel Xeon CPU, no GPU.
+For security and privacy reasons the data could not be move out of this environment.
+These constraints, commonly present in healthcare settings, influenced several pipeline design decisions, such as splitting the dataset into chunks for clustering, avoiding re-training Tranformer-based models and avoiding the use of generative large language models.
+2.3.2 Evaluation Procedure
+The pipeline was evaluated by comparing its output with the labels on the manually annotated subsets.
+Rather than applying the pipeline solely to the labelled records—which would hinder clustering performance due to limited data—we ran the full pipeline on the entire datasets and then extracted results for the annotated subsets.
+This ensured meaningful cluster formation and fair evaluation.
+We performed evaluation at two levels:
+  • Disease label evaluation: We assessed the accuracy of the disease labels produced by the pipeline by comparing them to ground-truth labels.
+To reduce noise, we retained only labels with more than 10 occurrences (1%) in the labelled subset, grouping all others into an Other category.
+Clusters not mapped to any label or unclustered clinical questions were assigned to the Other category.
+Metrics were computed at:
+    • Aggregate level: to measure overall precision, recall, and F1-score across all labels.
+    • Per-label level: to evaluate performance on individual categories.
+  • Appropriateness label evaluation: The quality of the pipeline’s assignment of appropriateness labels was evaluated using a four-class scheme: appropriate, inappropriate, likely (in)appropriate, and undetermined.
+The likely (in)appropriate category included referrals where reasons mentioned in the guidelines were present, but the text lacked complete information to assess the (in)appropriateness.
+The undetermined category was used for referrals not addressed in the clinical guidelines.
+Also in this case, we computed metrics and the aggregate and per-label levels.
+At both levels, we computed the metrics at the single clinical questions level and weighting clinical questions by their corresponding referral counts.
+The latter allows to incorporating the datasets’ real-world distributions into the analysis.
+We also qualitatively show t-SNE 2D projections of the embeddings of the labelled clinical questions.
+Although this type of visualization has known limitations [45], it remains a common approach in the literature for obtaining a qualitative overview of embedding structures [46].
+2.3.3 Hyperparameter Selection and Validation Strategy
+Clustering algorithm selection and hyperparameter tuning was carried out exclusively on the ECD dataset.
+The FEC dataset was reserved for external validation only, ensuring that all performance results reported on FEC represent generalization to unseen data and are free from overfitting to the development set.
+2.3.4 Comparative Evaluation
+To benchmark our pipeline, we compared it with alternative approaches using different text representations and/or clustering methods. These included:
+  • Embeddings from the original UmBERTo model (baseline)
+  • Embeddings from an Italian Word2Vec model [47]
+  • TF-IDF vector representations
+  • Latent Dirichlet Allocation (LDA)
+  • Direct matching between embeddings of clinical questions and embeddings of diseases from the guidelines
+Additional details are provided in Additional File 1, Section C.
+2.3.5 Regional and Temporal Analysis
+Finally, we applied the pipeline to the full datasets and reported results at a regional level.
+This provides a complete quantification of referrals for each identified category, together with their appropriateness.
+Moreover, we present an enrichment analysis stratifying results by year (pre-, during, and post-COVID pandemic), Local Health Authority (LHA), and prescribing physician type (general practitioners vs.
+specialists).
 3 Results
-Table 1: Results of clusters mapped to the labels defined by annotations on the manually labelled test set, computed considering clinical questions (CQ) and referrals (REF).
-P=Precision, R=Recall, F1=F1-Score.
-The values of all metrics are expressed as percentages.
+Table 1: Summary evaluation on the disease labels
 P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
+ECD
 Umberto-E3C-Pipeline 86.81 77.32 79.48 92.43 83.28 86.49
@@
 LDA 27.74 34.04 28.48 19.19 27.57 29.90
-Table 2: Results of the cluster mapping for each category, on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF).
-Labels with support <10absent10<10< 10 were grouped with other.
-P=Precision, R=Recall, F1=F1-Score.
+Embedding Match 68.29 58.44 59.21 70.59 61.52 62.09
+FEC
+Umberto-E3C-Pipeline 90.42 88.78 88.60 93.59 92.70 92.67
+Umberto-Base 84.56 75.24 79.63 88.58 79.66 83.88
+Word2Vec 81.21 70.55 75.51 82.45 73.86 77.92
+TF-IDF 58.42 51.23 54.59 60.41 53.49 56.74
+LDA 35.26 41.25 38.02 31.78 38.64 34.88
+Embedding Match 75.69 66.62 70.54 77.11 69.85 73.03
+00footnotetext: Summary results of mapping to the disease labels defined by manual annotations on the test sets, computed for clinical questions (CQ) and referrals (REF).
+P=Precision- R=Recall- F1=F1-Score.
+The values of all metrics are expressed as percentages.
+Table 2: Summary evaluation on the appropriateness labels
+P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
+ECD
+Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
+Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
+Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
+TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
+LDA 47.61 48.43 47.83 42.64 46.31 41.31
+Embedding Match 74.82 61.90 62.83 77.52 62.29 64.46
+FEC
+Umberto-E3C-Pipeline 92.05 90.96 90.88 94.66 93.96 94.05
+Umberto-Base 87.41 80.55 83.84 89.62 82.16 85.73
+Word2Vec 83.24 74.15 78.43 85.77 75.63 80.38
+TF-IDF 67.11 57.96 62.20 68.14 60.85 64.29
+LDA 38.96 45.62 42.03 35.41 42.93 38.81
+Embedding Match 80.07 69.92 74.20 81.15 73.92 77.15
+00footnotetext: Summary results of mapping to the appropriate/inappropriate/likely (in)appropriate/undetermined categories on the manually labelled test sets, computed for clinical questions (CQ) and referrals (REF).
+P=Precision- R=Recall- F1=F1-Score.
+The values of all metrics are expressed as percentages.
+3.1 Pipeline evaluation
+Table 1 presents the summary results of the evaluation performed on the manually annotated subsets, at the disease level, comparing different approaches.
+The pipeline with our fine-tuned Umberto embeddings demonstrates superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations.
+The approach with direct matching between embeddings of clinical questions and embeddings of diseases from guidelines exhibits discrete results, but lower than our pipeline, confirming that the clustering procedure act as a useful smoothing step.
+The metrics are computed for both clinical questions (CQ) and referrals (REF).
+The referral-level scores are higher for all methods, except LDA, due to the influence of clinical questions that occur more frequently, which seem to be easier to cluster for them, being less detailed.
+Table 3: Detailed results of the cluster mapping to the appropriateness labels for our pipeline
+P-CQ R-CQ F1-CQ #CQ P-REF R-REF F1-REF # REF
+ECD
+Appropriate 98.55 93.15 95.77 365 98.97 93.95 96.39 512
+Inappropriate 91.29 81.88 86.33 320 96.49 89.58 92.91 921
+Likely inappr. 43.28 89.23 58.29 65 84.03 89.89 90.86 628
+Undetermined 84.62 79.20 81.82 250 87.22 83.33 85.23 426
+FEC
+Appropriate 99.36 89.44 94.14 521 99.88 96.96 98.40 2501
+Inappropriate 95.35 87.23 91.11 47 95.71 88.16 91.78 76
+Likely appr. 98.27 87.18 92.39 195 98.88 90.72 94.62 388
+Undetermined 75.24 100.00 85.87 237 84.19 100.00 91.42 591
+00footnotetext: Detailed results of the cluster mapping to the appropriateness labels for our pipeline, on the manually labelled test sets, computed for clinical questions (CQ) and referrals (REF).
 P=Precision, R=Recall, F1=F1-Score.
 The values of all metrics are expressed as percentages.
+Table 2 shows the same results achieved by grouping labels into the appropriate, inappropriate, likely (in)appropriate and undetermined categories.
+These results show higher aggregate performance scores for ECD, compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories.
+The results for FEC are also superior, but the improvement is limited.
+This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging.
+Table 3 details the metric by appropriateness label, for our pipeline.
+At the clinical question level, it highlights issues with the likely inappropriate class for ECD, which corresponds directly to the varices class at disease level, and it shows slightly lower results for the undetermined class, but at the referrals level the only relevant difference is with the undetermined class for ECD.
+Table 4: Detailed results of the cluster mapping for each disease label for our pipeline on ECD dataset
 Category P-CQ R-CQ F1-CQ # CQ P-REF R-REF F1-REF # REF
 Claudicatio 83.33 45.45 58.82 11 90.00 56.25 69.23 16
-Pain 83.61 91.07 87.18 56 86.42 93.33 89.74 75
+Pain 85.00 91.07 87.93 56 88.61 93.33 90.91 75
 Edema 85.47 96.15 90.50 104 89.51 97.32 93.25 149
@@
 Ulcer 95.83 92.00 93.88 25 97.44 92.68 95.00 41
-Varices 41.43 89.23 56.59 65 82.47 98.89 89.83 628
+Varices 43.28 89.23 58.29 65 84.03 98.89 90.86 628
 Varices surgery 100.00 42.86 60.00 14 100.00 46.67 63.64 15
 Vasculopathy 82.14 44.23 57.50 52 95.33 68.92 80.00 148
-Other 65.37 67.94 65.68 131 62.56 70.59 66.33 187
-4.1 Pipeline evaluation
-Table 1 presents the results of the evaluation performed on the manually annotated subset.
-This evaluation compares the labels assigned to the clustered data with the ones defined by manual annotation.
-Clusters not mapped to any label or unclustered clinical questions were assigned to the other category.
-The pipeline demonstrated superior performance compared to alternative text representations and clustering algorithms, confirming the importance of contextual representations.
-The metrics were computed for both clinical questions (CQ) and referrals (REF).
-The referral-level scores were consistently higher due to the influence of clinical questions that occur more frequently, and that seem to be easier to cluster, being less detailed.
-Figure 4: Confusion matrix of the clustering assignment to the labels on the manually annotated test set, computed considering the number of clinical questions (A) and the number of referrals (B).
-[…]
-The confusion matrices shown in Figure 4 allow for a more detailed analysis of misclassification errors.
-Errors in varices surgery are often due to misclassifications as varices, a strictly related cluster.
-Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other.
-Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we noticed that some presented both diagnoses, making this misclassification reasonable.
-For the remaining classes, misclassified samples fall mainly into the other group, affecting its precision.
-Table 3: Results of clusters mapped to the appropriate/inappropriate/likely inappropriate/undetermined categories on the manually labelled test set, computed for clinical questions (CQ) and referrals (REF).
+Other 63.57 67.94 65.68 131 62.56 70.59 66.33 187
+00footnotetext: Detailed results of the cluster mapping for each disease for our pipeline, on the manually labelled test set for ECD, computed for clinical questions (CQ) and referrals (REF).
+Labels with support <10<10 are grouped with Other.
 P=Precision, R=Recall, F1=F1-Score.
 The values of all metrics are expressed as percentages.
-P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
-Umberto-E3C-Pipeline 89.24 88.27 88.64 93.58 91.52 92.51
-Umberto-Base 86.02 80.78 83.13 89.15 84.97 86.55
-Word2Vec 80.66 74.29 77.01 85.41 77.54 80.88
-TF-IDF 73.51 68.45 67.92 79.44 75.70 74.22
-LDA 47.61 48.43 47.83 42.64 46.31 41.31
-Table 3 shows the results achieved by grouping labels into the appropriate, inappropriate, likely inappropriate and undetermined categories.
-The results show higher aggregate performance scores compared to Table 1, indicating that most errors occurred within the same appropriateness category rather than across categories.
-This suggests that distinguishing between appropriate and inappropriate referrals is simpler, while finer distinctions at the label level are more challenging.
-A detailed list of all identified clusters, together with their mapping, is reported in Appendix B.
-Similar results, including validation on a manually annotated subset, are reported in Appendix D for another type of examination.
+Table 5: Detailed results of the cluster mapping for each disease label for our pipeline on FEC dataset
+Category P-CQ R-CQ F1-CQ # CQ P-REF R-RE F1-REF # REF
+Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
+Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
+Family history / prevention 97.56 86.33 91.60 139 98.57 91.39 94.85 302
+Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
+Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
+Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
+Chronic inflammatory bowel disease 100.00 89.29 94.34 56 100.00 88.37 93.83 86
+Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
+Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
+Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
+Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
+Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
+Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
+00footnotetext: Detailed results of the cluster mapping for each disease for our pipeline, on the manually labelled test set for FEC, computed for clinical questions (CQ) and referrals (REF).
+Labels with support <10<10 are grouped with Other.
+P=Precision, R=Recall, F1=F1-Score.
+The values of all metrics are expressed as percentages.
+Tables 4 and 5 detail metrics by disease for our pipeline.
+Most diseases achieve F1-CQ $80\%$, with many exceeding $90\%$.
+Lower scores are observed for some ECD categories such as varices, varices surgery, vasculopathy, claudicatio and other, while for FEC only other exhibits poor precision.
+At the referral level (F1-REF), only claudicatio, varices surgery and other achieve scores below $80\%$ for ECD, while for FEC only other is below $80\%$.
+Figure 5: Confusion matrix of the clustering assignment to the labels on the manually annotated test sets, for ECD (A, B) and FEC (C,D) computed considering the number of clinical questions (A, C) and the number of referrals (B, D)
+The confusion matrices shown in Figure 5 allow for a more detailed analysis of misclassification errors.
+For ECD, errors in varices surgery are often due to misclassifications as varices, a strictly related cluster.
+Regarding varices, its precision is mainly affected by misclassifications from venous insufficiency and other.
+Looking at the specific clinical questions manually labelled as venous insufficiency and classified as varices, we notice that some present both diagnoses, making this misclassification reasonable.
+For the remaining ECD classes, misclassified samples fall mainly into the other group, affecting its precision.
+For FEC, the majority of misclassification errors are wrong assignments to the other class.
+A few are wrong assignments to abdominal pain, which are in multiple cases due to the mention of this symptom along with others that were more relevant.
+A detailed list of all identified clusters, together with their mapping, is reported in Additional File 1, Section C.
+Figure 6 displays t-SNE 2D projections of the embedding vectors of the clinical questions in the two labelled subsets, showing that for most classes there are one or more distinct regions of the 2D space that tend to concentrate the corresponding clinical questions.
+Figure 6: 2D t-SNE projections of embeddings of clinical questions from the validation sets of ECD (A) and FEC (B) with their labels, excluding the Other class
 3.2 Analysis of the entire dataset
-Table 4: Summary of referrals’ clusters, mapped to all labels defined in manual annotations.
-Green indicates appropriate clusters, red inappropriate clusters, yellow clusters more likely to be inappropriate, and light grey those not mentioned in the guidelines.
+Table 6: Summary of referrals’ clusters for ECD with appropriateness labels
 Cluster Label # REF %
-Venous insufficiency 102,339 20,63
-Thrombosis 98,821 19,92
-Varices 71,269 14,37
-Edema 61,822 12,46
-(thrombo)phlebitis 49,537 9,99
-Pain 21,131 4,26
-Vasculopathy 11,244 2,27
-Saphenectomy 8,290 1,67
-Varices surgery 7,126 1,44
-Diabetes 6,583 1,33
-Ulcer 6,272 1,26
-Paresthesias 4,856 0,98
-Claudicatio 3,570 0,72
-Dimer 1,601 0,32
-Swelling 1,238 0,25
-Respiratory insufficiency 1,060 0,21
-Heavy legs 705 0,14
-Ischaemia 617 0,12
-Stent 612 0,12
-Dermatitis 474 0,10
-Hypercolesterolaemia 416 0,08
-Pregnancy 225 0,05
-Dialisis 222 0,04
-Surgery 198 0,04
-Smoker 193 0,04
-Dyspnoea 171 0,03
-Tingling 170 0,03
-Cold legs 147 0,03
-Hypoesthesia 80 0,02
-Metabolic syndrome 48 0,01
-Bleeding 29 0,01
-Other 34,917 7,04
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+00footnotetext: Summary of referrals’ clusters for ECD, mapped to all labels defined in manual annotations.
+Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely inappropriate, and light grey those not mentioned in the guidelines.
+Table 7: Summary of referrals’ clusters for FEC with appropriateness labels
+Cluster Label # REF %
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+00footnotetext: Summary of referrals’ clusters for FEC, mapped to all labels defined in manual annotations.
+Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
+Tables 3.2 and 1 extend the analysis to the entire ECD and FEC datasets, highlighting the appropriateness labels for each cluster.
+We can observe that the distribution of the reasons for the referrals is unbalanced, with 75% of referrals covered by the first five reasons in ECD and by the first seven reasons in FEC.
+In ECD, varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to exclude, even for a human reader.
+In FEC, colon cancer familiarity / prevention and chronic inflammatory bowel disease were instead marked as likely appropriate since guidelines state that physicians should carefully consider the patient history and condition to establish if and when they should be performed.
+Specific conditions mentioned in the guidelines — such as requiring ten years from diagnosis or significant colon involvement — were largely absent from referral texts.
+Summarizing Table 3.2, for ECD, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate.
+The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations.
+For FEC, 55.98% of referrals appear to be appropriate, while only 5.33% seems to be not appropriate.
+There are then 15.95% more likely to be appropriate, 15.77% that are not mentioned on the guidelines, and 6.98% that cannot be clustered or cannot be mapped to the labels defined by manual annotations.
+Notably, in both datasets the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines, especially for ECD.
+Although individually rare, these uncovered reasons accounted for a non-negligible proportion of overall referrals, suggesting the need for a broader guideline scope.
+Additionally, we observed ambiguity in referrals citing varices, where physicians often omitted whether surgical intervention was required.
+In contrast, FEC referrals showed higher levels of appropriateness, though some inappropriate clusters were identified.
+Figure 7 provides some examples, along with their English translation, of clustered clinical questions.
 Figure 7: Examples of clustered clinical questions. English translations do not completely reflect the abbreviations and typos of the original Italian texts.
-Table 4 extends the analysis on the entire dataset, highlighting the appropriateness labels for each cluster.
-It can be observed that the distribution of the reasons for the referrals is highly unbalanced, with the first five covering more than 75% of referrals.
-Varices cluster was assigned to likely inappropriate since this diagnosis is inappropriate unless it is associated with surgery, an element that is often difficult to completely exclude, even for a human reader.
-Summarizing the table, 34.32% of referrals appear to be appropriate, while 34.07% seems to be not appropriate.
-The remaining 31.61% referrals can be further divided into 14.37% more likely to be not appropriate, 10.20% that are not mentioned on the guidelines, and 7.04% that cannot be clustered or cannot be mapped to the labels defined by manual annotations.
-Notably, the other group is limited, while there are clusters for many symptoms and diseases not mentioned in the guidelines.
-Figure 5 provides some examples, along with their English translation, of clustered clinical questions.
-Figure 6: Prescription appropriateness grouped by LHA (A), type of prescribing physicians (B) and year (C)
+Figure 8: Prescription appropriateness stratified by LHA (A, B), type of prescribing physicians (C, D) and year (E, F) for ECD (A,C,E) and FEC (B,D,F).
 Figure 8 reports the level of appropriateness stratified by LHA, type of prescribing physicians and year, providing additional insights into the reasons for inappropriate prescriptions.
+This stratified analysis reveals significant geographical variability for both examinations and especially for ECD, suggesting that targeted interventions are needed in specific Local Health Authorities.
+General practitioners exhibited higher rates of inappropriate ECD referrals compared to specialist physicians.
+This might have multiple reasons, including a larger patient base, more frequent encounters with less severe cases, and a lower knowledge and/or agreement with guidelines [48, 49].
+For FEC, the rate of inappropriateness is similar but specialized physicians have more appropriate referrals.
+An important trend emerges when analyzing the impact of COVID-19 on referral practices.
+During the early pandemic phase in 2020, inappropriate ECD referrals dropped significantly, likely due to resource constraints prompting physicians to prioritize urgent cases.
+This pattern is consistent with existing literature, which reports a general reduction in non-urgent medical procedures and a shift toward more conservative referral behaviour during the pandemic [50, 51, 52, 53].
+However, this decline was temporary, with inappropriate referrals rebounding in 2021 as healthcare services normalized.
+A similar, though less pronounced, trend was observed for FEC, possibly due to the already limited number of inappropriate referrals for this examination.
 4 Discussion
-In this study, we analyzed texts of Italian referrals for prescription appropriateness, with a case study on venous echocolordoppler of lower limbs.
-By introducing a novel approach that clusters referral texts and maps these clusters to existing guidelines, we demonstrated how it is possible to extract meaningful and actionable insights into the reasons behind such prescriptions.
-Our work has several strengths.
-The analysis encompassed 496,971 referrals, covering an entire Italian region with 10 million inhabitants for three years.
-This large-scale dataset enabled a comprehensive evaluation of prescription practices.
-Validation was performed on a manually annotated subset, demonstrating that contextual embedding models significantly outperformed simpler clustering techniques in capturing the semantic nuances of referral texts.
-Analyzing the results in detail, we observed very high performances for most diseases/symptoms, with only two groups with lower recall.
-Additionally, results on a secondary dataset (reported in Appendix D) confirmed the generalizability of our methodology.
-The analysis of the whole dataset allowed us to have, for the first time, a clear picture of referral motivations for this examination.
-This is a relevant result for public health authorities, considering the absence of structured recording of this information in the Italian NHS.
-Our mapping of referrals to guidelines revealed frequent inappropriate uses, as well as numerous reasons not explicitly covered by current guidelines.
-While individually rare, these reasons cumulatively represent a non-negligible share of referrals, suggesting the need to expand the scope of guidelines.
-The ambiguity in referrals grouped under varices, where physicians often omit whether surgery is indicated, together with generic reasons like pain and vasculopathy, also points to the importance of requesting more specific information from referring physicians.
-Stratification revealed significant geographical variability, suggesting that targeted interventions are needed in specific Local Health Authorities.
-General practitioners exhibited higher rates of inappropriate referrals compared to specialists.
-This might have multiple reasons, including larger patient base, more frequent encounter with less severe cases, but also a lower knowledge and/or agreement with guidelines [45, 46].
-Temporal trends showed a decline in inappropriate referrals during the initial COVID-19 wave in 2020, followed by an increase in 2021 as the healthcare system normalized.
-This aligns with literature suggesting that during the pandemic’s peak, physicians prioritized urgent and necessary examinations to manage resource constraints effectively [47, 48, 49, 50].
+In this study, we proposed a methodological pipeline leveraging NLP techniques to assess the appropriateness of diagnostic referrals in the Italian NHS.
+We illstrated its use through two case studies on venous echocolordoppler of the lower limbs in angiology and flexible endoscope colonoscopy in gastroenterology.
+By applying an unsupervised approach, we uncovered key trends in referral motivations and guideline adherence, offering the first large-scale assessment of this kind in Italy.
+Our work presents several strengths.
+The pipeline operates in an unsupervised setting, without the need for labelled data, which are costly and time-consuming to produce.
+This makes it applicable to any examination with established guidelines defined in terms of appropriate and inappropriate reasons for referrals.
+It is also designed to run on large datasets without requiring high computational resources, lowering the barrier for adoption.
+In addition to quantifying appropriateness, the pipeline offers a detailed picture of referral motivations, thereby revealing reasons not explicitly covered by existing guidelines.
+We demonstrated these capabilities on two population-scale datasets, covering an entire Italian region with 10 million inhabitants for three years.
+Validation on manually annotated subsets confirmed the robustness of our pipeline and demonstrated that contextual embedding models significantly outperform simpler clustering techniques in capturing the semantic nuances of referral texts.
+We observed consistently high performances at the appropriate/inappropriate level and for most diseases/symptoms, with only two clusters in ECD with lower recall and one cluster in FEC with lower precision.
+The excellent results on two different examinations, one of which was not used during pipeline development, confirmed the generalizability of our methodology.
+The large-scale analysis enabled, for the first time, a comprehensive view of referral motivations for these examinations.
+This is a relevant result for public health authorities, given the absence of structured recording of this information in the Italian NHS.
+The identification of undocumented reasons for referrals highlighted potential gaps in the existing guidelines and the stratified analyses further revealed geographical and temporal patterns and variations across physician types, consistent with findings in the literature and suggestive of targeted interventions to reduce inappropriate referrals.
+Importantly, the findings directly supported the Lombardy Region in issuing a new resolution extending and reinforcing referral guidelines.
+This represents one of the first documented cases in Italy where an AI-based analysis of routine healthcare data directly shaped health policy.
+This study is a relevant example of a real-world, large-scale application of artificial intelligence in healthcare.
+Despite a growing body of research, the adoption of AI-based solutions in clinical practice faces significant challenges [54] and remains largely dominated by diagnostic tools in medical imaging [55, 56], particularly in Italy [57].
+Our work points to a complementary direction: leveraging large-scale textual data to monitor referral practices.
+In this way, it illustrates how AI can extend its role beyond diagnosis to support healthcare governance and inform evidence-based policy.
 Despite its strengths, this study has some limitations.
-Manual annotation was performed on a subset of data due to the large time required, and annotations were conducted by a single annotator, preventing inter-annotator agreement analysis.
-While valid for most cases, the focus on a single reason per referral overlooks referrals with multiple reasons, which could be addressed through fuzzy clustering.
-Another limitation was the lack of access to patient history, which prevented a precise assessment of the appropriateness of certain referrals, such as those for varices, where the presence of a surgery often could not be verified.
-It is also essential to note that this analysis does not aim to state the appropriateness of single referrals, for which a comprehensive analysis of the clinical status and history of the patient would be necessary.
-Instead, it aims to provide a comprehensive, automatized and accurate picture at the regional level.
-In future research, we aim to expand the validation of our pipeline to other types of examinations, with larger manually annotated datasets, ensuring broader applicability across healthcare services.
+Manual annotation was limited to a subset of the data due to time constraints, and performed by a single annotator, limiting the possibility of inter-annotator agreement analysis.
+Our approach also assigns a single primary reason per referral, potentially overlooking multi-reason referrals, which could be better addressed through fuzzy clustering techniques.
+Moreover, the lack of access to patient histories limited our ability to fully assess clinical appropriateness in certain cases, such as referrals for varices, where prior surgical indication is relevant.
+It is important to clarify that our pipeline does not aim to assess the appropriateness of individual referrals.
+Instead, our approach provides a systematic, automated, and large-scale evaluation of prescription practices at the regional level, offering valuable insights for policymakers and healthcare administrators.
+In future work, we aim to extend the validation of the pipeline to other examinations and broader settings, including larger annotated datasets and other Italian regions, to further support the adoption of data-driven tools for evaluating referral practices and improving healthcare delivery.
 5 Conclusions
-This study proposed a two-step procedure for assessing the level of appropriateness of Italian referrals.
-We presented a case study on the venous echocolordoppler of the lower limbs in the Lombardy Region, demonstrating excellent performance on a validation set and extracting relevant insights for public health authorities.
-The outcomes of this analysis enabled the Lombardy Region to reinforce existing guidelines for physicians, aiming to reduce the number of inappropriate referrals.
-By addressing one of the key drivers of inefficiency in the healthcare system, this approach has the potential to alleviate significant burdens on the Italian NHS, lowering waiting times and enhancing the overall effectiveness of healthcare services.
+This study proposed a methodological pipeline for assessing the level of appropriateness of Italian referrals.
+We presented two case studies on venous echocolordoppler of the lower limbs and flexible endoscope colonoscopy in the Lombardy Region, demonstrating excellent performance on manually annotated test sets and deriving relevant insights for public health authorities.
+The outcome of this analysis allowed the Lombardy Region to reinforce existing guidelines for physicians, with the aim of reducing the number of inappropriate referrals.
+This work is an example of how artificial intelligence can be deployed on large, real-world administrative datasets, highlighting the potential of AI to support healthcare systems in monitoring practices, informing policy, and ultimately alleviating significant burdens on the Italian NHS by lowering waiting times and enhancing overall effectiveness of care.
+Format: pdf Title: Supplementary Material Description: contains more details about the manual annotation process, hyperparameters selection and additional results
 […]
@@
   1. 1.
+    AI: Artificial Intelligence
+  2. 2.
+    BERT: Bidirectional Encoder Representation for Transformers
+  3. 3.
+    CIBD: Chronic Inflammatory Bowel Disease
+  4. 4.
+    CQ: Clinical Question
+  5. 5.
+    DGR: Decreto della Giunta Regionale (Regional Law)
+  6. 6.
+    ECD: venous EcoColorDoppler of the lower limbs
+  7. 7.
+    EHR: Electronic Health Record
+  8. 8.
+    FEC: Flexible Endoscope Colonscopy with possible biopsies
+  9. 9.
+    GP: General Practitioner
+  10. 10.
+    ICD: International Classification of Diseases
+  11. 11.
+    IQR: InterQuartile Range
+  12. 12.
+    LDA: Latent Dirichlet Allocation
+  13. 13.
+    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
+  14. 14.
     NHS: National Health Service
-  2. 2.
+  15. 15.
     NLP: Natural Language Processing
-  3. 3.
-    GP: General Practitioner
-  4. 4.
-    EHR: Electronic Health Record
-  5. 5.
-    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
-  6. 6.
-    GDPR: General Data Protection Regulation
-  7. 7.
-    IQR: Interquartile Range
-  8. 8.
-    ICD: International Classification of Diseases
-  9. 9.
-    BERT: Bidirectional Encoder Representation for Transformers
-  10. 10.
-    DGR: Decreto della Giunta Regionale (Regional Law)
-  11. 11.
+  16. 16.
     REF: Referral
-  12. 12.
-    CQ: Clinical Question
+  17. 17.
+    TF-IDF: Term Frequency - Inverse Document Frequency
 Declarations
@@
 4 c.1 (https://gdpr.eu/article-4-definitions/) and Recital 26 (https://gdpr-info.eu/recitals/no-26/).
+Consent for publication
+Not applicable.
 Availability of data and materials
@@
 Code availability
-The code is available by the authors upon request.
 Competing interests
 The authors declare that they have no competing interests.
+Funding
+No fundings supported this research.
 Authors’ contributions
-VT: investigation, formal analysis, methodology, software, data curation, writing - original draft AB: resources, data curation, conceptualization ME: resources, data curation, conceptualization OL: resources, data curation, conceptualization FI: conceptualization, methodology, formal analysis, supervision, writing - review & editing
-Acknowledgments
+VT: analysis and interpretation of data, development of the methodology, drafting of the manuscript AB: concept and design of the study, acquisition, analysis and interpretation of data ME: concept and design of the study, acquisition, analysis and interpretation of data OL: concept and design of the study, acquisition, analysis and interpretation of data FI: concept and design of the study, analysis and interpretation of data, critical revision of the manuscript, supervision.
+All authors read and approved the final manuscript.
+Acknowledgements
 The present research is part of the activities of “Dipartimento di Eccellenza 2023-2027”.
@@
 Appendix A Manual annotations
-A validation set of 1000 clinical questions was manually annotated for the main disease/symptom determining the referrals.
-The existing guidelines from the Lombardy Region [26] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate.
+A validation set of 1,000 clinical questions was manually annotated for the main disease/symptom determining the referrals, for both ECD and FEC datasets.
+The existing guidelines from the Lombardy Region [31, 32] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate.
 If none of the categories mentioned in the guidelines fit the reason for the referral, a new category representing the main disease or symptom mentioned in the referral was created.
@@
     If there are multiple symptoms/diseases not mentioned in the guidelines, prefer those for which a category was already created, otherwise the first one to be mentioned in the text
-This procedure, after four iterations, including consultation with an angiologist, defined the labels reported in Table A.1.
-Table A.1: Labels derived by manual annotations on the validation set of 1000 clinical questions for venous echocolordoppler of the lower limbs. CQ = Clinical Questions, REF = Referrals.
+This procedure, after four iterations, including consultation with specialist physicians, defined the labels reported in Tables A.1 and A.2. After review of these labels, some of them were merged:
+  • bleeding and occult blood since bleeding in this context is always an occult bleeding
+  • colitis, chron’s disease, ulcerative rectocolitis, chronic inflammatory bowel disease, under the latter which is a category comprehensive of all these diseases
+Table A.1: Labels derived by manual annotations on the validation set of 1,000 clinical questions for ECD.
 Label # CQ # REF
@@
 Vascular encephalopathy 1 1
-Appendix B Clustering
-B.1 Additional details on clustering procedures
-For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens.
-Clustering was performed with HDBscan and K-Means, obtaining the best results with K-Means.
-The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40.
-For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7.
-For HDBscan, the initial minimum size was set to 500, selected as the minimum size that led to the emergence of at least 5 clusters, and was progressively halved up to 20.
-B.2 Additional clustering results
-Table LABEL:tab:clusters-summ presents all the clusters identified for venous echocolordoppler of the lower limbs, with their keyword descriptions, the sizes and the labels they were assigned.
-Table B.1: Summary of all clusters identified for echocolordoppler of the lower limbs, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
-CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
-claudicatio; riferita claudication; referred claudication 1132 1686
-claudicatio; intermittens claudicatio; intermittens claudication 653 1179
-claudicatio claudicatio claudicatio 402 530
-claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudicatio 60 126
-claudicatio; arti; inferiori claudicatio; limbs; lower claudicatio 41 49
-dermatite dermatitis dermatitis 421 474
-diabete diabetes diabetes 1540 3249
-diabete; mellito diabetes; mellitus diabetes 534 1847
-diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
-diabetico; diabetica diabetes; diabetic diabetes 209 165
-dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
-dialisi dialysis dialysis 110 222
-dimero dimer dimer 1343 1601
-dispnea dyspnoea dyspnoea 104 171
-dolore pain pain 6343 12051
-deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
-algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
-polpaccio; dolore calf; pain pain 707 1753
-algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
-dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
-dolore; popliteo pain; popliteal pain 35 117
-algia; covid; gamba pain; covid; leg pain 30 63
-dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
-crampi; notturni cramps; nocturnal pain 6 5
-edema edema edema 11686 20723
-edemi; declivi edema; declivities edema 9491 16473
-edema; arti; edemi edema; limbs; edemas edema 3712 7867
-arti; inferiori; edemi limbs; lower; edemas edema 528 2486
-edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
-edema; arti; inferiori edema; limbs; lower edema 899 2350
-edema; perimalleolare edema; perimalleolar edema 1039 2214
-edema; natura; determinare edema; nature; determine edema 1027 2199
-edemi; perimalleolari edema; perimalleolar edema 1190 2151
-edema; malleolare edema; malleolar edema 935 1136
-edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
-edema; flebite edema; phlebitis edema 217 297
-arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
-arti; inferiori; edema limbs; lower; edema edema 80 168
-arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
-edema; inferiori; arti edema; lower; limbs edema 74 163
-edema; sinistra edema; left edema 61 108
-claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
-linfedema; arti lymphedema; limbs edema 23 24
-emorragia haemorrhage haemorrhage 18 29
-formicolio tingling tingling 43 170
-freddo cold cold 49 147
-fumatore smoker smoker 121 193
-gravidanza pregnancy pregnancy 92 225
-insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
-venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
-venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
-superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
-insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
-venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
-venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
-venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
-insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
-venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
-venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
-insufficienza; venosa venous; insufficiency venous insufficiency 0 0
-arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
-venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
-venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
-venosa; cronica venous; chronic venous insufficiency 0 0
-insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
-intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
-lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
-lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
-ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
-ipoestesia hypoesthesia hypoesthesia 31 80
-ischemia ischaemia ischaemia 517 617
-parestesie paresthesias paresthesias 1834 3965
-parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
-pesantezza heaviness heaviness 215 475
-arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
-arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
-safenectomia saphenectomy saphenectomy 1988 4177
-safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
-safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
-safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
-sindrome metabolica metabolic syndrome metabolic syndrome 12 48
-stent stent stent 301 612
-tromboflebite thrombophlebitis thrombophlebitis 7073 17415
-flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
-flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
-varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
-tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
-gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
-tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
-flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
-flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
-tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
-flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
-tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
-trombosi thrombosis thrombosis 17060 43558
-trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
-trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
-trombosi; venosa venous; thrombosis thrombosis 1325 3337
-venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
-poplitea; trombosi popliteal; thrombosis thrombosis 607 642
-venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
-trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
-segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
-safena; trombosi saphenous vein; thrombosis thrombosis 301 263
-post; trombotica post; thrombotic thrombosis 36 73
-tumefazione swelling swelling 533 741
-tumefazione; gamba swelling; leg swelling 172 448
-destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
-ulcera; vascolare ulcer; vascular ulcer 828 3861
-ulcera ulcer ulcer 1258 2411
-varici varices varices 13999 35939
-arti; inferiori; varici limbs; lower; varices varices 7018 18157
-venose; varici venous; varices varices 4957 13554
-varicosita; stasi varicosities; stasis varices 957 2701
-varici; arti; inferiori varices; limbs; lower limbs varices 219 277
-vene; varicose; algie veins; varicose; algie varices 121 194
-safena; varici saphenous veins; varicose varices 47 136
-varicosita; varicosita’ varicosities; varicositis varices 86 97
-vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
-varici; sospette varicose veins; suspected varices 34 63
-varicosita; gamba varicosities; leg varices 80 63
-varici; intervento varicosities; surgery varices intervention 3144 7126
-vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
-vasculopatia vasculopathy; peripheral vasculopathy 769 1886
-vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
-arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
-arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
-iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
-aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
-vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
-vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
-vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
-arteriopatia arteriopathy vasculopathy 20 25
-natura; determinare; periferica nature; determine; peripheral other 3978 10039
-natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
-paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
-arti; inferiori limbs; lower other 2190 5165
-safena; incontinenza saphenous; incontinence other 288 782
-ipertrofia hypertrophy other 506 591
-tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
-embolia; polmonare embolism; pulmonary other 110 174
-femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
-stenosi; venoso stenosis; venous other 30 49
-dislipidemia; ipertensione dyslipidaemia; hypertension other 25 44
-aorta; sottorenale aorta; subrenal other 25 29
-Table B.1: *
-Appendix C Pre-processing results
-Table C.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table C.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing.
-Table C.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
+Table A.2: Labels derived by manual annotations on the validation set of 1,000 clinical questions for FEC
+Label # CQ # REF
+Polyps 158 344
+Bleeding 127 959
+Colon cancer 107 187
+Abdominalgia 104 214
+Colon cancer familiarity 82 141
+Anaemia 65 115
+Prevention 57 161
+Occult blood 36 789
+Constipation 36 139
+Diverticulosis 33 48
+Altered bowel 30 65
+Colitis 28 40
+Chronic diarrhoea 28 107
+Diverticulitis 14 28
+Crohn’s disease 12 16
+Haemorrhoidal disease 9 30
+Chronic inflammatory bowel disease 8 15
+Ulcerative rectocolitis 8 15
+Surgery 5 6
+Intestinal obstruction 5 10
+Dyspepsia 4 5
+Weight loss 3 8
+Irritable bowel syndrome 3 6
+Lymphoma 3 4
+Transplant 3 15
+Other 2 2
+Colonpathy 2 7
+Enteritis 2 3
+Proctitis 2 4
+Rectal prolapse 2 12
+Lynch syndrome 2 2
+Angiodysplasia 1 1
+Appendicitis 1 1
+Celiac disease 1 1
+Intestinal loops distension 1 1
+Hepatopathy 1 1
+Protruding formation 1 1
+Gastritis 1 2
+Ileitis 1 1
+Hypercholesterolaemia 1 1
+Splenic flexure thickening 1 3
+Rectal wall thickening 1 2
+Sigma wall thickening 1 1
+Behcet’s disease 1 1
+Liver disease 1 1
+Pancreatitis 1 1
+Proctalgia 1 3
+Sigmoiditis 1 1
+Suspected colic lesion 1 33
+Splenic fissure stenosis 1 1
+Rectal tenesmus 1 2
+Appendix B Pre-processing results
+Table B.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table B.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing.
+Table B.3 reports the summary results on disease labels without applying pre-processing.
+Table B.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
 Abbr Freq Expansion Expansion [ENG]
@@
 $t$ 203 tibio-tarsica tibio-tarsal
-Table C.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.
+Table B.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.
 Original Word Corrected Word Corrected Word [ENG]
@@
 acugene acufene tinnitus
-Appendix D Additional results on gastroenterology exam
-Additional results related to the flexible endoscope colonoscopy with possible biopsies examination are reported in this section.
-The dataset for this exam consisted of all referrals filled in in the Lombardy Region between 2019 and 2021, for a total of 407,949 referrals.
-A random sample of 1,000 distinct clinical questions (corresponding to 3,556 referrals) was annotated in the same way as for the case study presented in the main manuscript and used to validate the results.
-Considering 2015 guidelines from Regione Lombardia for this type of exam [27], we marked colon cancer familiarity/prevention and chronic inflammatory bowel disease as likely appropriate.
-Guidelines state that, in these cases, physicians should carefully consider the patient history and condition to establish if and when the exam should be performed.
-Table D.1 reports the detailed results per category, while Table D.2 the summary metrics considering the categories and the appropriate/inappropriate levels.
-Table D.3 presents the final results on the full dataset, which are stratified by year, type of prescribing physician and LHU in Figure D.1.
-Table D.4 details the labels from manual annotation on the test set.
-Table D.1: Detailed results of the cluster mapping for each category, on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF).
+Table B.3: Results of clusters mapped to the disease labels on the manually labelled test sets, computed for clinical question (CQ) and referrals (REF) without applying pre-processing steps.
 P=Precision, R=Recall, F1=F1-Score.
-Category P-CQ [%] R-CQ [%] F1-CQ [%] # CQ P-REF [%] R-REF [%] F1-REF[%] # REF
-Bleeding 100.00 77.91 87.59 163 100.00 97.31 98.64 1748
-Polyps 99.35 97.47 98.40 158 99.71 98.55 99.12 344
-Family history / prevention 97.56 86.33 91.60 139 98.57 91.39 94.85 302
-Colon cancer 96.84 85.98 91.09 107 98.24 89.30 93.56 187
-Abdominal pain 88.57 89.42 89.00 104 93.66 89.72 91.65 214
-Anaemia 92.75 98.46 95.52 65 95.73 97.39 96.55 115
-Chronic inflammatory bowel disease 100.00 89.29 94.34 56 100.00 88.37 93.83 86
-Constipation 78.38 80.56 79.45 36 91.61 94.24 92.91 139
-Diverticulosis 90.00 81.82 85.71 33 86.67 81.25 83.87 48
-Altered bowel 84.85 93.33 88.89 30 92.54 95.38 93.94 65
-Chronic diarrhoea 100.00 82.14 90.20 28 100.00 93.46 96.62 107
-Diverticulitis 100.00 92.86 96.30 14 100.00 89.29 94.34 28
-Other 47.14 98.51 63.77 67 59.93 99.42 74.78 173
-Table D.2: Summary metrics on the manually labelled test set for flexible endoscope colonscopy, computed for clinical questions (CQ) and referrals (REF), considering annotations categories or only appropriate/inappropriate/likely appropriate/undetermined.
-P=Precision, R=Recall, F1=F1-Score.
-P-CQ [%] R-CQ [%] F1-CQ [%] P-REF [%] R-REF [%] F1-REF [%]
-Annotation categories 90.42 88.78 88.60 93.59 92.70 92.67
-Appropriateness levels 92.05 90.96 90.88 94.66 93.96 94.05
-Table D.3: Summary of referrals’ clusters for flexible endoscope colonscopy, mapped to all labels defined in manual annotations.
-Green indicates appropriate clusters, red inappropriate clusters, yellow clusters for which the information might not be sufficient to define the appropriateness but is more likely appropriate, and light grey those not mentioned in the guidelines.
-Cluster Label # REF %
-polyps 81,111 19.88
-occult blood 76,487 18.75
-colon cancer familiarity / prevention 43,615 10.69
-abdominalgia 31,316 7.68
-colon cancer 30,434 7.46
-anaemia 23,133 5.67
-chronic inflammatory bowel disease 21,470 5.26
-constipation 15,109 3.70
-diverticulosis 15,058 3.69
-chronic diarrhoea 13,533 3.32
-altered alvo 10,933 2.68
-haemorrhoidal disease 5,379 1.32
-diverticulitis 4,669 1.14
-suspected colic lesion detected by other investigation technique 3,671 0.90
-irritable bowel syndrome 2,023 0.50
-surgery 797 0.20
-lynch syndrome 787 0.19
-other 28,424 6.98
-Figure D.1: Prescription appropriateness for flexible endoscope colonscopy grouped by LHA (A), type of prescribing physicians (B) and year (C)
-Table D.4: Labels derived by manual annotations on the validation set of 1000 clinical questions for flexible endoscope colonoscopy.
-Label # CQ # REF
-polyps 158 344
-bleeding 122 954
-colon cancer 107 187
-abdominalgia 104 214
-colon cancer familiarity 82 141
-anaemia 65 115
-prevention 57 161
-occult blood 36 789
-constipation 36 139
-diverticulosis 33 48
-altered bowel 30 65
-colitis 28 40
-chronic diarrhoea 28 107
-diverticulitis 14 28
-crohn’s disease 12 16
-haemorrhoidal disease 9 30
-chronic inflammatory bowel disease 8 15
-ulcerative rectocolitis 8 15
-surgery 5 6
-intestinal obstruction 5 10
-proctorrhagia 5 5
-dyspepsia 4 5
-weight loss 3 8
-irritable bowel syndrome 3 6
-lymphoma 3 4
-transplant 3 15
-other 2 2
-colonpathy 2 7
-enteritis 2 3
-proctitis 2 4
-rectal prolapse 2 12
-lynch syndrome 2 2
-angiodysplasia 1 1
-appendicitis 1 1
-celiac disease 1 1
-intestinal loops distension 1 1
-hepatopathy 1 1
-protruding formation 1 1
-gastritis 1 2
-ileitis 1 1
-hypercholesterolaemia 1 1
-splenic flexure thickening 1 3
-rectal wall thickening 1 2
-sigma wall thickening 1 1
-behcet’s disease 1 1
-liver disease 1 1
-pancreatitis 1 1
-proctalgia 1 3
-sigmoiditis 1 1
-suspected lesion 1 33
-splenic fissure stenosis 1 1
-rectal tenesmus 1 2
+The values of all metrics are expressed as percentages.
+P-CQ R-CQ F1-CQ P-REF R-REF F1-REF
+ECD
+Umberto-E3C-Pipeline 84.01 73.51 78.40 89.15 79.28 83.95
+Umberto-Base 77.99 70.68 74.18 85.01 73.18 78.66
+Word2Vec 75.63 62.25 68.27 78.90 67.77 72.99
+TF-IDF 46.28 33.61 38.92 45.73 37.83 41.46
+LDA 24.38 32.27 27.88 16.51 25.97 20.24
+FEC
+Umberto-E3C-Pipeline 88.59 85.53 87.12 90.99 87.65 89.32
+Umberto-Base 80.28 70.31 75.02 85.79 75.42 80.35
+Word2Vec 78.66 68.55 73.27 80.15 70.38 74.96
+TF-IDF 55.45 44.91 49.69 55.95 49.60 52.65
+LDA 31.52 36.30 33.80 26.70 33.91 29.86
+Appendix C Clustering
+C.1 Hyperparameters and algorithms selection
+Table LABEL:tab:hdbscan_hyper reports the grid search for HDBSCAN hyperparameters, evaluated with F1-CQ on ECD dataset.
+Table C.1: Grid search for HDBSCAN hyperparameters, with F1-CQ on ECD dataset
+min_cluster_size min_samples cluster_selection_epsilon F1-CQ [%]
+20 0 11.42
+20 0.1 11.42
+20 0.5 11.42
+20 1 0 63.91
+20 1 0.1 63.91
+20 1 0.5 63.91
+20 5 0 10.13
+20 5 0.1 10.13
+20 5 0.5 10.13
+20 10 0 77.85
+20 10 0.1 77.85
+20 10 0.5 77.85
+20 20 0 11.42
+20 20 0.1 11.42
+20 20 0.5 11.42
+50 0 40.31
+50 0.1 40.31
+50 0.5 40.31
+50 1 0 60.81
+50 1 0.1 60.81
+50 1 0.5 60.81
+50 5 0 71.72
+50 5 0.1 71.72
+50 5 0.5 71.72
+50 10 0 70.40
+50 10 0.1 70.40
+50 10 0.5 70.40
+50 20 0 40.31
+50 20 0.1 40.31
+50 20 0.5 40.31
+100 0 23.87
+100 0.1 23.87
+100 0.5 23.87
+100 1 0 75.64
+100 1 0.1 75.64
+100 1 0.5 75.64
+100 5 0 44.60
+100 5 0.1 44.60
+100 5 0.5 44.60
+100 10 0 51.82
+100 10 0.1 51.82
+100 10 0.5 51.82
+100 20 0 40.31
+100 20 0.1 40.31
+100 20 0.5 40.31
+200 0 14.46
+200 0.1 14.46
+200 0.5 14.46
+200 1 0 67.41
+200 1 0.1 67.41
+200 1 0.5 67.41
+200 5 0 36.43
+200 5 0.1 36.43
+200 5 0.5 36.43
+200 10 0 43.32
+200 10 0.1 43.32
+200 10 0.5 43.32
+200 20 0 31.80
+200 20 0.1 31.80
+200 20 0.5 31.80
+500 0 18.64
+500 0.1 18.64
+500 0.5 18.64
+500 1 0 15.48
+500 1 0.1 15.48
+500 1 0.5 15.48
+500 5 0 14.45
+500 5 0.1 14.45
+500 5 0.5 14.45
+500 10 0 23.37
+500 10 0.1 23.37
+500 10 0.5 23.37
+500 20 0 16.31
+500 20 0.1 16.31
+500 20 0.5 16.31
+step from 500 0 34.52
+step from 500 0.1 34.52
+step from 500 0.5 34.52
+step from 500 1 0 71.54
+step from 500 1 0.1 71.54
+step from 500 1 0.5 71.54
+step from 500 5 0 79.48
+step from 500 5 0.1 79.48
+step from 500 5 0.5 79.48
+step from 500 10 0 63.28
+step from 500 10 0.1 63.28
+step from 500 10 0.5 63.28
+step from 500 20 0 41.95
+step from 500 20 0.1 41.95
+step from 500 20 0.5 41.95
+Table C.1: *
+Figures C.1 and C.2 report the performance of K-Means and Agglomerative Hierarchical Clustering on the ECD dataset, at the CQ level, varying the number of clusters.
+Their results are consistently lower than those from HDBSCAN.
+Figure C.1: K-means performance on ECD-CQ varying the number of clusters
+Figure C.2: Agglomerative hierarchical clustering performance on ECD-CQ varying the number of clusters
+For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens.
+The best results presented are obtained with K-Means.
+The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40 for ECD and k=35 for FEC.
+For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7 for ECD and k=9 for FEC.
+For direct embeddings match, we matched Umberto embeddings of clinical questions with embeddings of known diseases/symptoms from guidelines, using cosine similarity.
+If no element in the guidelines matched with a similarity of at least 0.6 we assigned the Other label.
+C.2 Additional clustering results
+Tables LABEL:tab:clusters-summ and LABEL:tab:clusters-summ-fec present all the clusters identified for ECD and FEC, respectively, with their keyword descriptions, the sizes and the labels they were assigned.
+Table C.2: Summary of all clusters identified for ECD, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
+ID CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
+1 claudicatio; riferita claudication; referred claudication 1132 1686
+2 claudicatio; intermittens claudicatio; intermittens claudication 653 1179
+3 claudicatio claudicatio claudication 402 530
+4 claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudication 60 126
+5 claudicatio; arti; inferiori claudicatio; limbs; lower claudication 41 49
+6 dermatite dermatitis dermatitis 421 474
+7 diabete diabetes diabetes 1540 3249
+8 diabete; mellito diabetes; mellitus diabetes 534 1847
+9 diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
+10 diabetico; diabetica diabetes; diabetic diabetes 209 165
+11 dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
+12 dialisi dialysis dialysis 110 222
+13 dimero dimer dimer 1343 1601
+14 dispnea dyspnoea dyspnoea 104 171
+15 dolore pain pain 6343 12051
+16 deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
+17 algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
+18 polpaccio; dolore calf; pain pain 707 1753
+19 algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
+20 dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
+21 dolore; popliteo pain; popliteal pain 35 117
+22 algia; covid; gamba pain; covid; leg pain 30 63
+23 dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
+24 crampi; notturni cramps; nocturnal pain 6 5
+25 edema edema edema 11686 20723
+26 edemi; declivi edema; declivities edema 9491 16473
+27 edema; arti; edemi edema; limbs; edemas edema 3712 7867
+28 arti; inferiori; edemi limbs; lower; edemas edema 528 2486
+29 edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
+30 edema; arti; inferiori edema; limbs; lower edema 899 2350
+31 edema; perimalleolare edema; perimalleolar edema 1039 2214
+32 edema; natura; determinare edema; nature; determine edema 1027 2199
+33 edemi; perimalleolari edema; perimalleolar edema 1190 2151
+34 edema; malleolare edema; malleolar edema 935 1136
+35 edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
+36 edema; flebite edema; phlebitis edema 217 297
+37 arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
+38 arti; inferiori; edema limbs; lower; edema edema 80 168
+39 arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
+40 edema; inferiori; arti edema; lower; limbs edema 74 163
+41 edema; sinistra edema; left edema 61 108
+42 claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
+43 linfedema; arti lymphedema; limbs edema 23 24
+44 emorragia haemorrhage haemorrhage 18 29
+45 formicolio tingling tingling 43 170
+46 freddo cold cold 49 147
+47 fumatore smoker smoker 121 193
+48 gravidanza pregnancy pregnancy 92 225
+49 insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
+50 venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
+51 venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
+52 superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
+53 insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
+54 venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
+55 venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
+56 venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
+57 insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
+58 venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
+59 venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
+60 insufficienza; venosa venous; insufficiency venous insufficiency 0 0
+61 arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
+62 venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
+63 venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
+64 venosa; cronica venous; chronic venous insufficiency 0 0
+65 insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
+66 intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
+67 lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
+68 lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
+69 ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
+70 ipoestesia hypoesthesia hypoesthesia 31 80
+71 ischemia ischaemia ischaemia 517 617
+72 parestesie paresthesias paresthesias 1834 3965
+73 parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
+74 pesantezza heaviness heaviness 215 475
+75 arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
+76 arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
+77 safenectomia saphenectomy saphenectomy 1988 4177
+78 safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
+79 safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
+80 safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
+81 sindrome metabolica metabolic syndrome metabolic syndrome 12 48
+82 stent stent stent 301 612
+83 tromboflebite thrombophlebitis thrombophlebitis 7073 17415
+84 flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
+85 flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
+86 varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
+87 tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
+88 gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
+89 tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
+90 flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
+91 flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
+92 tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
+93 flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
+94 tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
+95 trombosi thrombosis thrombosis 17060 43558
+96 trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
+97 trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
+98 trombosi; venosa venous; thrombosis thrombosis 1325 3337
+99 venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
+100 poplitea; trombosi popliteal; thrombosis thrombosis 607 642
+101 venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
+102 trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
+103 segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
+104 safena; trombosi saphenous vein; thrombosis thrombosis 301 263
+105 post; trombotica post; thrombotic thrombosis 36 73
+106 tumefazione swelling swelling 533 741
+107 tumefazione; gamba swelling; leg swelling 172 448
+108 destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
+109 ulcera; vascolare ulcer; vascular ulcer 828 3861
+110 ulcera ulcer ulcer 1258 2411
+111 varici varices varices 13999 35939
+112 arti; inferiori; varici limbs; lower; varices varices 7018 18157
+113 venose; varici venous; varices varices 4957 13554
+114 varicosita; stasi varicosities; stasis varices 957 2701
+115 varici; arti; inferiori varices; limbs; lower limbs varices 219 277
+116 vene; varicose; algie veins; varicose; algie varices 121 194
+117 safena; varici saphenous veins; varicose varices 47 136
+118 varicosita; varicosita’ varicosities; varicositis varices 86 97
+119 vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
+120 varici; sospette varicose veins; suspected varices 34 63
+121 varicosita; gamba varicosities; leg varices 80 63
+122 varici; intervento varicosities; surgery varices intervention 3144 7126
+123 vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
+124 vasculopatia vasculopathy; peripheral vasculopathy 769 1886
+125 vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
+126 arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
+127 arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
+128 iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
+129 aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
+130 vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
+131 vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
+132 vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
+133 arteriopatia arteriopathy vasculopathy 20 25
+134 natura; determinare; periferica nature; determine; peripheral other 3978 10039
+135 natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
+136 paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
+137 arti; inferiori limbs; lower other 2190 5165
+138 safena; incontinenza saphenous; incontinence other 288 782
+139 ipertrofia hypertrophy hypertrophy 506 591
+140 tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
+141 embolia; polmonare embolism; pulmonary embolism 110 174
+142 femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
+143 stenosi; venoso stenosis; venous other 30 49
+144 dislipidemia; ipertensione dyslipidaemia; hypertension dyslipidaemia 25 44
+145 aorta; sottorenale aorta; subrenal other 25 29
+Table C.2: *
+Table C.3: Summary of all clusters identified for FEC, including their keyword representation, the label assigned to them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
+ID CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
+1 addominalgia; alvo abdominalgia; alvo abdominal pain 163 371
+2 addominalgie; natura; determinare abdominalgia; nature; determine abdominal pain 714 5.801
+3 addominali; algie; dolori abdominalgia; algia; pain abdominal pain 894 5.430
+4 addominali; coliche abdominal; colic abdominal pain 240 3.775
+5 alvo; addominalgie bowel; abdominal pain abdominal pain 224 2.507
+6 calo; ponderale; addominalgia fall; weight; abdominalgia abdominal pain 100 215
+7 natura; determinare; coliche nature; determine; colic abdominal pain 325 2.590
+8 natura; determinare; addominali; coliche nature; determine; abdominal; colic abdominal pain 3.687 5.605
+9 fianco; dolore; persistente flank; pain; persistent abdominal pain 61 121
+10 fossa; addominalgia fossa; abdominalgia abdominal pain 192 2.125
+11 fossa; iliaca; algia fossa; iliac; algia abdominal pain 90 194
+12 fossa; iliaca; algie fossa; iliac; algia abdominal pain 33 71
+13 fossa; iliaca; dolore fossa; iliac; pain abdominal pain 302 662
+14 ipocondrio; destra; dolore; sinistra hypochondrium; right; pain; left abdominal pain 127 281
+15 quadranti; addominalgia; addominalgie quadrant; abdominalgia; abdominalgia abdominal pain 33 74
+16 stipsi; addominalgia constipation; abdominalgia abdominal pain 501 1.105
+17 stipsi; addominali constipation; abdominal abdominal pain 112 250
+18 stipsi; addominali; algie constipation; abdominal; pain abdominal pain 61 139
+19 alvo; alterazione; alterazioni alvo; alteration altered bowel 432 3.500
+20 alvo; alterno bowel; abdominalgia altered bowel 610 2.496
+21 alvo; alterno; addominalgia bowel; abdominalgia altered bowel 287 1.925
+22 alvo; calo; ponderale bowel; decline; weight altered bowel 196 1.655
+23 alvo; stipsi bowel; constipation altered bowel 129 277
+24 fossa; alvo fossa; alvo altered bowel 197 1.080
+25 anemia; emoglobina anaemia; haemoglobin anaemia 61 128
+26 anemia; natura; determinare; grave anaemia; nature; determine; severe anaemia 602 4.392
+27 anemia; sangue; occulto; positivo anaemia; blood; occult; positive anaemia 315 2.366
+28 anemia; sideropenica; origine; indeterminata anaemia; sideropenic; origin; undetermined anaemia 332 6.663
+29 anemia; sideropenica; sangue; occulto anaemia; sideropenic; blood; occult anaemia 49 112
+30 anemizzazione; grave; severa; sangue anaemia; severe; blood anaemia 456 3.814
+31 anemizzazione; natura; determinare anaemia; nature; determine anaemia 250 2.504
+32 sideropenica; anemia iron deficiency; anaemia anaemia 365 3.154
+33 diarrea; colite; cronica diarrhoea; colitis; chronic chronic diarrhoea 95 757
+34 diarrea; cronica diarrhoea; chronic chronic diarrhoea 778 7.668
+35 diarroico; alvo diarrhoeic; alvo chronic diarrhoea 193 2.694
+36 natura; determinare; diarrea; dolore; addominale nature; determine; diarrhoea; pain; abdominal chronic diarrhoea 409 2.414
+37 colite; ischemica colitis; ischaemic chronic inflammatory bowel disease 119 286
+38 crohn; morbo crohn disease; disease chronic inflammatory bowel disease 540 3.812
+39 crohn; morbo; ileo crohn; disease; ileum chronic inflammatory bowel disease 71 202
+40 infiammatoria; malattia; intestinale; cronica inflammatory; disease; intestinal; chronic chronic inflammatory bowel disease 220 2.333
+41 malattie; infiammatorie; croniche; intestinali diseases; inflammatory; chronic; intestinal; disease chronic inflammatory bowel disease 313 3.780
+42 malattie; infiammatorie; croniche; intestinali; calprotectina inflammatory; disease; chronic; intestinal; calprotectin chronic inflammatory bowel disease 19 51
+43 stipsi; colite constipation; colitis chronic inflammatory bowel disease 74 185
+44 ulcerosa; colite ulcerative; colitis chronic inflammatory bowel disease 2.985 7.323
+45 ulcerosa; rettocolite ulcerative; rectocolitis chronic inflammatory bowel disease 283 3.498
+46 adenocarcinoma; colon adenocarcinoma; colon colon cancer 336 753
+47 adenoma; adenomi; colon adenoma; adenomas; colon colon cancer 50 102
+48 alvo; tumore; colon bowel; cancer; colon colon cancer 57 120
+49 anamnesi; colon anamnesis; colon colon cancer 76 301
+50 colico; tumore colic; cancer colon cancer 97 219
+51 colon; adenoma; adenomi; linfoma colon; adenoma; lymphoma colon cancer 123 394
+52 colon; discendente; tumore colon; descending; cancer colon cancer 61 137
+53 colon; pregressa; asportazione; tumore colon; previous; removal; cancer colon cancer 41 80
+54 colon; ricerca; resezione; dopo; sincrone; cancro; lesioni; curativa colon; research; resection; after; synchronous; cancer; lesions; curative colon cancer 171 2.878
+55 ematochezia; tumore; colon haematochezia; cancer; colon colon cancer 389 867
+56 emicolectomia; tumore; esiti; destra; colon haemolectomy; cancer; outcomes; right; colon colon cancer 452 1.055
+57 neoformazione; rettale; retto; colon neoformation; rectal; rectum; colon colon cancer 51 119
+58 neoplasia; colon neoplasia; colon colon cancer 438 1.323
+59 neoplasia; eteroplasia neoplasm; heteroplasm colon cancer 113 289
+60 sigma; tumore sigma; cancer colon cancer 197 449
+61 stipsi; tumore; colon constipation; cancer; colon colon cancer 286 682
+62 tumore; colon cancer; colon colon cancer 5.725 16.757
+63 tumore; colon; controllo cancer; colon; control colon cancer 108 324
+64 tumore; colon; famigliarita cancer; colon; familial colon cancer 346 1.111
+65 tumore; intestinale cancer; intestine colon cancer 305 840
+66 tumore; pregresso; colon cancer; proctorragia; colon colon cancer 306 815
+67 tumore; proctorragia; colon cancer; proctorragia; colon colon cancer 71 148
+68 tumore; retto cancer; rectum colon cancer 266 671
+69 ostinata; stipsi obstinate; constipation constipation 2.356 5.962
+70 stipsi; cronica constipation; chronic constipation 745 7.521
+71 stipsi; fossa; iliaca constipation; iliac fossa constipation 27 59
+72 stipsi; ostinata; addominali constipation; obstinate; abdominal constipation 88 190
+73 stipsi; ponderale; calo constipation; ponderal; fall constipation 122 1.377
+74 diverticolite diverticulitis diverticulitis 759 3.148
+75 diverticolite; acuta; recente diverticulitis; acute; recent diverticulitis 373 1.172
+76 sigma; diverticolite sigma; diverticulitis diverticulitis 101 221
+77 sigmoidite; sospetta; recente; episodio; diverticolare; esiti; diverticolite; diverticolosi sigmoiditis; suspected; recent; episode; diverticular; outcome; diverticulitis; diverticulosis diverticulitis 55 128
+78 alvo; diverticolosi bowel; diverticulosis diverticulosis 359 2.126
+79 diverticolare; malattia diverticular; disease diverticulosis 147 438
+80 diverticoli; diverticolare; colon; poliposi diverticula; diverticular; colon; polyposis diverticulosis 388 1.849
+81 diverticolosi; addominalgia diverticulosis; abdominalgia diverticulosis 355 782
+82 diverticolosi; addominali diverticulosis; abdominal diverticulosis 264 2.213
+83 diverticolosi; colon diverticulosis; colon diverticulosis 337 3.281
+84 diverticolosi; colon; diverticolite diverticulosis; colon; diverticulitis diverticulosis 253 754
+85 diverticolosi; sigma diverticulosis; sigma diverticulosis 117 256
+86 pregressa; asportazione; diverticolosi previous; removal; diverticulosis diverticulosis 412 2.096
+87 pregressa; polipectomia; diverticolosi previous; polypectomy; diverticulosis diverticulosis 142 306
+88 stipsi; diverticolosi constipation; diverticulosis diverticulosis 136 957
+89 emorroidaria; sindrome; patologia haemorrhoidal; syndrome; pathology haemorrhoidal disease 113 259
+90 prevenzione; emorroidi cancer prevention / familiarity; haemorrhoids haemorrhoidal disease 962 4.647
+91 prolasso; emorroidario prolapse; haemorrhoids haemorrhoidal disease 182 473
+92 lynch; sindrome; colon lynch; syndrome; colon lynch syndrome 140 787
+93 addominalgia; proctorragia abdominalgia; proctorragia bleeding 99 217
+94 alvo; ematochezia bowel; haematochezia bleeding 116 251
+95 alvo; sangue; occulto; fecale bowel; blood; occult; faecal bleeding 175 382
+96 alvo; sanguinamento; irregolare; digestivo bowel; bleeding; irregular; digestive bleeding 703 4.915
+97 ematiche; perdite bleeding; leakage bleeding 151 639
+98 ematochezia; positivo; sanguinamento; natura; determinare haematochezia; positive; bleeding; nature; determine bleeding 760 12.945
+99 ematochezia; addominalgia; addominalgie haematochezia; abdominalgia; abdominalgia bleeding 61 134
+100 ematochezia; ematochezie haematochezia; haematochezia bleeding 500 5.307
+101 melena; riferita; episodio melaena; reported; episode bleeding 182 953
+102 nuovi; anticoagulanti; orali; terapia; proctorragia; anemia; paziente new; anticoagulants; oral; therapy; proctorragia; anaemia; patient bleeding 66 152
+103 occulto; alvo; sangue; positivo occult; alvo; blood; positive bleeding 41 86
+104 occulto; fecale; sangue; anemia occult; faecal; blood; anaemia bleeding 824 1.917
+105 occulto; sangue occult; blood bleeding 4.459 12.755
+106 occulto; sangue; positivo occult; blood; positive bleeding 33 75
+107 oncologica; prevenzione oncological; cancer prevention / familiarity bleeding 2.346 4.752
+108 proctorragia proctorrhagia bleeding 279 1.236
+109 proctorragia; diverticolosi proctorrhagia; diverticulosis bleeding 152 328
+110 proctorragia; emorroidi; episodi proctorrhagia; haemorrhoids; episodes bleeding 571 1.439
+111 proctorragia; rettorragia proctorragia; rectorrhagia bleeding 4.219 13.288
+112 proctorragie; emorroidi proctorragia; haemorrhoids bleeding 65 220
+113 prolasso; proctorragia; emorroidario prolapse; proctorragia; haemorrhoid bleeding 46 98
+114 rettorragia; colon rectorrhagia; colon bleeding 48 104
+115 rettorragia; episodi rectorrhagia; episodes bleeding 408 1.167
+116 rettorragia; tumore; colon rectorrhagia; cancer; colon bleeding 74 157
+117 sangue; episodi; presenza blood; episodes; presence bleeding 330 1.647
+118 sangue; occulto blood; occult bleeding 263 619
+119 sangue; occulto; anemizzazione; positivo blood; occult; anaemia; positive bleeding 122 292
+120 sangue; occulto; calprotectina; fecale; positivo blood; occult; calprotectin; faecal; positive bleeding 25 64
+121 sangue; occulto; fecale blood; occult; faecal bleeding 2.525 7.258
+122 sangue; occulto; fecale; anemizzazione blood; occult; faecal; anaemia bleeding 224 494
+123 sangue; occulto; fecale; calo; ponderale blood; fecal; occult; weight-loss bleeding 140 312
+124 sangue; occulto; fecale; positivita blood; fecal; occult; positive bleeding 198 529
+125 sangue; occulto; fecale; positivo; terapia; anemia blood; faecal; occult; positive; therapy; anaemia bleeding 38 96
+126 sangue; occulto; fecale; tumore blood; occult; faecal; cancer bleeding 171 364
+127 sangue; occulto; polipectomia blood; occult; polypectomy bleeding 105 250
+128 sangue; occulto; positivo blood; occult; positive bleeding 69 151
+129 sangue; stipsi; occulto; fecale blood; constipation; occult; faecal bleeding 163 357
+130 stipsi; ematochezia constipation; haematochezia bleeding 91 190
+131 stipsi; proctorragia constipation; proctorragia bleeding 156 347
+132 adenomatosi; polipi adenomatous; polyps polyps 49 116
+133 asportato; colon; polipo removed; colon; polyp polyps 182 424
+134 asportazione; polipo removal; polyp polyps 457 1.052
+135 colon; polipi colon; polyps polyps 2.079 6.733
+136 exeresi; polipi; pregressa exeresis; polyps; previous polyps 71 174
+137 follow; up; polipectomia follow; up; polypectomy polyps 133 333
+138 polipectomia; colon polypectomy; colon polyps 89 236
+139 polipectomia; controllo polypectomy; control polyps 516 2.263
+140 polipectomia; endoscopica; pregressa polypectomy; endoscopic; previous polyps 109 356
+141 polipectomia; pregressa; addominalgia polypectomy; previous; abdominalgia polyps 49 111
+142 polipectomia; sigma polypectomy; sigma polyps 80 177
+143 polipectomia; tumore; colon polypectomy; cancer; colon polyps 32 65
+144 polipi; adenomi; sorveglianza; pazienti; serrati polyps; adenomas; surveillance; patients; serratus polyps 932 23.390
+145 polipi; colon polyps; colon polyps 610 15.287
+146 polipi; controllo polyps; control polyps 365 4.950
+147 polipi; intestinali polyps; intestinal polyps 186 881
+148 polipi; pregressi polyps; previous polyps 333 1.128
+149 polipo; adenomatoso polyps; adenomatous polyps 44 104
+150 polipo; anale polyp; anal polyps 123 340
+151 polipo; colon polyp; colon polyps 168 5.634
+152 poliposi; colon polyposis; colon polyps 711 1.847
+153 poliposi; intestinale polyposis; intestinal polyps 231 1.876
+154 poliposi; pregressa polyposis; previous polyps 53 119
+155 poliposi; sospetta polyposis; suspected polyps 263 875
+156 pregressa; polipectomia previous; polypectomy polyps 981 6.133
+157 pregressa; poliposi previous; polyposis polyps 412 2.208
+158 pregressa; sigma; asportazione; polipo previous; sigma; removal; polyp polyps 79 165
+159 pregresso; polipo previous; polyp polyps 322 3.204
+160 proctorragia; poliposi; colon proctorragia; polyposis; colon polyps 98 228
+161 sigma; polipo sigma; polyp polyps 272 621
+162 stipsi; ostinata; poliposi constipation; obstinate; polyposis polyps 38 81
+163 alvo; familiarita; tumore alvo; familial; cancer cancer prevention / familiarity 212 604
+164 anamnesi; tumore; colon; familiare anamnesis; cancer; colon; familial cancer prevention / familiarity 200 3.358
+165 colon; familiarita; tumore colon; familial; cancer cancer prevention / familiarity 302 6.530
+166 diagnosi; precoce; prestaz; diagnost; tumori early; diagnosis; performance; diagnostics; cancers cancer prevention / familiarity 106 1.378
+167 familiarita; tumore; colon familial; cancer; colon cancer prevention / familiarity 1.137 3.286
+168 famigliarita; tumore; colon familial; cancer; colon cancer prevention / familiarity 151 388
+169 familiarita; colon; neoplasia familiarity; colon; neoplasm cancer prevention / familiarity 1.160 3.450
+170 familiarita; prevenzione familiarity; cancer prevention / familiarity cancer prevention / familiarity 306 1.114
+171 oncologica; familiarita; prevenzione; tumore oncological; familial; cancer prevention / familiarity; cancer cancer prevention / familiarity 195 453
+172 prestaz; diagnost; diagnosi; precoce; tumori prior; diagnosis; early; diagnosis; cancers cancer prevention / familiarity 55 136
+173 prestaz; diagnost; diagnosi; precoce; tumori; tumore; familiarita; colon service; diagnosis; early; diagnosis; cancers; cancer; familial; colon cancer prevention / familiarity 192 418
+174 prestaz; diagnost; diagnosi; sangue; occulto; precoce; tumori performance; diagnosis; diagnosis; occult; blood; early; cancers cancer prevention / familiarity 49 126
+175 tumore; anamnesi; familiare cancer; history; family history cancer prevention / familiarity 280 3.603
+176 tumore; colon; familiarita cancer; colon; familial cancer prevention / familiarity 618 4.379
+177 tumore; colon; familiarita’ cancer; colon; familial cancer prevention / familiarity 401 5.010
+178 tumore; colon; proctorragia; familiarita cancer; colon; proctorship; familiarity cancer prevention / familiarity 127 274
+179 tumore; familiarita cancer; familiarity cancer prevention / familiarity 3.709 8.696
+180 tumore; intestino; familiarita cancer; intestine; familiarity cancer prevention / familiarity 64 412
+181 asportazione; adenoma; esiti removal; adenoma; outcome surgery 45 91
+182 esiti; resezione; retto outcomes; resection; rectum surgery 57 47
+183 pregressa; asportazione previous; removal surgery 921 105
+184 pregressa; asportazione; adenoma previous; removal; adenoma surgery 180 293
+185 pregressa; emicolectomia previous; hemicolectomy surgery 75 85
+186 renale; trapianto; lista; insufficienza; cronica; dialisi renal; transplantation; list; failure; chronic; dialysis surgery 25 64
+187 trapianto; fegato transplant; liver surgery 86 112
+188 colica; sospetta; lesione; riscontrata; altra; tecnica; indagine colic; suspected; lesion; found; other; technique; investigation suspected colic lesion 179 3.671
+189 calo; ponderale fall; weight other 333 1.022
+190 colon; irritabile colon; irritable other 263 1.432
+191 displasia; adenoma; tubulare; grado dysplasia; adenoma; tubular; grade other 148 316
+192 fossa; iliaca fossa; iliac other 24 54
+193 grado; displasia grade; dysplasia other 84 191
+194 prolasso; rettale prolapse; rectal other 57 122
+195 sigma; ispessimento sigma; thickening other 43 87
+196 subocclusione; intestinale sub-occlusion; intestinal other 81 202